### v3

# A uniform-in-time nonlocal approximation of the standard Fokker-Planck equation

Abstract
We study a nonlocal approximation of the Fokker-Planck equation in which we can estimate the speed of convergence to equilibrium in a way which does not degenerate as we approach the local limit of the equation. This uniform estimate cannot be easily obtained with standard inequalities or entropy methods, but can be obtained through the use of Harris’s theorem, finding interesting links to quantitative versions of the central limit theorem in probability. As a consequence one can prove that solutions of this nonlocal approximation converge to solutions of the usual Fokker-Planck equation uniformly in time—hence we show the approximation is asymptotic-preserving in this sense. The associated equilibrium has some interesting tail and regularity properties, which we also study.
1 Introduction
We consider the linear partial differential equation
$\partial_{t}u=\frac{1}{\varepsilon^{2}}\big(J_{\varepsilon}*u-u\big)+\operatorname{div}(xu)=:L_{\varepsilon}u,$ (NLFP)
where $u\colon[0,+\infty)\times\mathbb{R}^{d}\to\mathbb{R}$ is the unknown whose variables we usually name by $u\equiv u(t,x)$, and $\varepsilon>0$ is a scaling parameter. This PDE involves a nonnegative kernel $J\colon\mathbb{R}^{d}\to[0,+\infty)$ which we always assume at least integrable, and such that
$\begin{split}\int_{\mathbb{R}^{d}}J(x)\,\mathrm{d}x&=1,\qquad\int_{\mathbb{R}^{d}}J(x)x\,\mathrm{d}x=0\qquad\int_{\mathbb{R}^{d}}J(x)x_{i}x_{j}\,\mathrm{d}x=2\delta_{ij},\end{split}$ (1)
for $i,j=1,\dots,d$. Making straightforward scaling arguments one can also apply our results if $J$ is not of mass $1$.
A canonical example of $J$ for which our results are meaningful is any smooth, compactly supported probability distribution satisfying the above, which is a helpful case to keep in mind. The scaled kernel $J_{\varepsilon}$ is defined by
$J_{\varepsilon}(x):=\varepsilon^{-d}J\Big(\frac{x}{\varepsilon}\Big),\qquad x\in\mathbb{R}^{d},$
and the convolution in (NLFP) is understood to be in the $x$ variable. Thus (NLFP) is, strictly speaking, a partial integro-differential equation since it involves an integral kernel. We choose this particular scaling in $\varepsilon$ since it is well known that as $\varepsilon\to 0$ the PDE (NLFP) resembles the standard Fokker-Planck PDE
$\partial_{t}u=\Delta u+\operatorname{div}(xu)=:L_{0}u.$ (FP)
We refer to the book by Andreu-Vaillo et al. (2010) for a good explanation of this limit. Hence, equation (NLFP) can be understood as a nonlocal version of the standard Fokker-Planck equation.
Our interest in studying equation (NLFP) is not directly due to the model itself, but as a proof of concept of a new method to study nonlocal-to-local limits. Equation (NLFP) is a nonlocal diffusion which is interesting enough to have non-explicit equilibria, which likely do not have Gaussian decay (see Section 5.3). Standard entropy methods are hard to apply to this PDE, especially in the limit $\varepsilon\to 0$. Though the equation can be studied by Fourier methods (see Andreu-Vaillo et al. (2010)), we are interested in methods which do not depend on an explicit Fourier transform solution. This said, we also point out that the drift $\operatorname{div}(xu)$ appears in a natural way if we consider the PDE $\partial_{t}v=J*v-v$ and then carry out the standard diffusive change of variables $u(t,x)=e^{dt}v(\frac{1}{2}e^{2t},xe^{t})$. The equation for $u$ then becomes
$\partial_{t}u=\frac{1}{\varepsilon(t)^{2}}\left(J_{\varepsilon(t)}*u-u\right)+\operatorname{div}(xu),$ (2)
where now $\varepsilon(t)=e^{-t}$. Equation (NLFP) is then the result of “freezing” the parameter $\varepsilon$ in this equation, and as such is a natural model to understand regarding nonlocal diffusion.
In contrast with local diffusion, the nonlocal PDE (NLFP) does not have instantaneous regularization properties. Both PDE (NLFP) and (FP) have a stochastic interpretation in terms of a Markov process; we refer to standard texts such as Risken (1996) for this. In this paper we are interested in understanding the asymptotic behavior of solutions to (NLFP) as $t\to+\infty$ and as $\varepsilon\to 0$ by the use of a new method based on Harris’s theorem, since techniques for nonlocal PDE still have many shortcomings and this equation is a good test case. We pay particular attention to whether the asymptotic behavior of (NLFP) matches that of the standard Fokker-Planck equation as $\varepsilon\to 0$. The use of Harris’s theorem is by no means new in the study of nonlocal PDE, but as far as we know it is the first time that it can be applied to give a result which is stable under the scaling in (NLFP). We consider this the main novelty of this paper.
Obviously, since our results will be valid also for a fixed $\varepsilon$, we may frame the problem in a slightly different way by studying the PDE
$\partial_{t}u=c\big(J*u-u\big)+\operatorname{div}(xu),$
for $J$ a general probability distribution satisfying (1). We have chosen to write the PDE as in (NLFP) in order to emphasize that we are especially interested in its scaling limit to the Fokker-Planck equation.
Summary of main results
Here are the two main results we will prove in this paper:
  1. 1.
    Solutions to (NLFP) converge to equilibrium exponentially:
    $\|u(t,\cdot)-F_{\varepsilon}\|\leq Ce^{-\gamma t}\|u_{0}-F_{\varepsilon}\|,\qquad t\geq 0$
    in a variety of weighted $L^{1}$ norms $\|\cdot\|$, where $F_{\varepsilon}$ is a suitable equilibrium of the PDE. Our main point is that both constants $C$ and $\gamma$ can be taken to be independent of $\varepsilon$ as $\varepsilon\to 0$. This is not contained in previous results, except for Mischler and Tristani (2017), who use different techniques.
  2. 2.
    If $u$ is the solution to the nonlocal equation (NLFP) with a symmetric $J$ and $v$ the solution to the standard Fokker-Planck equation (FP) with the same initial condition, then
    $\|u(t,\cdot)-v(t,\cdot)\|\leq C\varepsilon^{2}\qquad\text{for all $t\geq 0$,}$
    again in a variety of weighted $L^{1}$ norms. The constant $C$ is independent of $\varepsilon$ and $t$, so the convergence happens uniformly for all times. We refer to this as “asymptotic-preserving convergence”, and is not proved in previous literature as far as we know.
Let us describe these results more precisely. We always denote by $L_{1}^{k}$ the usual weighted Lebesgue space with norm
$\left\|\varphi\right\|_{L^{1}_{k}}=\int_{\mathbb{R}^{d}}(1+|x|^{k})|\varphi(x)|\,\mathrm{d}x$
or with the equivalent norm
$\left\|\varphi\right\|_{L^{1}_{\langle k\rangle}}=\int_{\mathbb{R}^{d}}\left\langle x\right\rangle^{k}|\varphi(x)|\,\mathrm{d}x,$
where $\left\langle x\right\rangle^{k}:=(1+|x|^{2})^{k/2}$, a “smoothed $k$-th power” which avoids regularity problems at $x=0$. (We denote by $\left\langle x\right\rangle$ the smoothed $1^{st}$ power.) Similarly we define $L^{1}_{\exp,a}$ as the Lebesgue space with weight $e^{a|x|}$ with natural norm
$\left\|\varphi\right\|_{\exp,a}=\int_{\mathbb{R}^{d}}e^{a|x|}\left|\varphi(x)\right|\,\mathrm{d}x$
or with equivalent norm
$\left\|\varphi\right\|_{\left\langle\exp\right\rangle,a}=\int_{\mathbb{R}^{d}}e^{a\left\langle x\right\rangle}\left|\varphi(x)\right|\,\mathrm{d}x.$
An equilibrium of equation (NLFP) is a function $F_{\varepsilon}\in L^{1}$ which satisfies the stationary PDE
$\frac{1}{\varepsilon^{2}}\big(J_{\varepsilon}*F_{\varepsilon}-F_{\varepsilon}\big)+\operatorname{div}(xF_{\varepsilon})=0$
in the distributional sense. A probability equilibrium is a nonnegative equilibrium $F_{\varepsilon}$ with $\int_{\mathbb{R}^{d}}F_{\varepsilon}(x)\,\mathrm{d}x=1$. It is easy to see that there exists at most one probability equilibrium (see Theorem (11)) and under mild additional conditions we also show existence (and additional estimates). Here is the main theorem we prove in this paper regarding asymptotic behavior of mild solutions.
Theorem 1.1.
Let $\varepsilon\in(0,1]$ and let $J\colon\mathbb{R}^{d}\to[0,+\infty)$ satisfy (1) and, moreover, $J\in L^{p}$ for some $p\in(1,\infty]$. Then,
  1. (i)
    If $J\in L^{1}_{k^{\prime}}$ for some $k^{\prime}>2$, then there exists a unique equilibrium $F_{\varepsilon}$ of (NLFP) in $L^{1}_{k^{\prime}}$. Additionally, for every $0<k\leq k^{\prime}$ there exist (explicit) constants $C\geq 1$ and $\gamma>0$, independent of $\varepsilon$, such that any (mild) solution $u$ of (NLFP) with a probability initial condition $u_{0}\in L^{1}_{k}$ satisfies
    $\left\|u(t,\cdot)-F_{\varepsilon}\right\|_{L^{1}_{k}}\leq Ce^{-\gamma t}\left\|u_{0}-F_{\varepsilon}\right\|_{L^{1}_{k}}\qquad\text{for every $t\geq 0$}.$
  2. (ii)
    If $J\in L^{1}_{\exp,a^{\prime}}$, for some $a^{\prime}>0$, then there exists a unique equilibrium $F_{\varepsilon}$ of (NLFP) in $L^{1}_{\exp,a}$, for all $a<a^{\prime}$. Additionally, there exist (explicit) constants $C\geq 1$ and $\gamma>0$, independent of $\varepsilon$, such that any (mild) solution $u$ of (NLFP) with a probability initial condition $u_{0}\in L^{1}_{\exp,a}$ satisfies
    $\left\|u(t,\cdot)-F_{\varepsilon}\right\|_{L^{1}_{\exp,a}}\leq Ce^{-\gamma t}\left\|u_{0}-F_{\varepsilon}\right\|_{L^{1}_{\exp,a}}\qquad\text{for every $t\geq 0$}.$
    We remark that the unique equilibrium $F_{\varepsilon}$ also belongs to the (smaller, but $\varepsilon$-dependent) space $L^{1}_{\exp,b}$ for any $b<a^{\prime}\varepsilon^{-1}$.
The dependence of the constants can be explicitly traced throughout the proof. Among the various factors involved, the constants in general get better the closer $J$ resembles a Gaussian, in a sense that will be clarified later.
Remark 1.2.
From our proof of the theorem (or directly as a simple consequence) we get the same estimates substituting $u-F_{\varepsilon}$ by any solution $w$ with initial mass 0 (i.e. $\int_{\mathbb{R}^{d}}w_{0}(x)\,\mathrm{d}x=0$). For example, we prove that under the conditions of point (i),
$\left\|w(t,\cdot)\right\|_{L^{1}_{k}}\leq Ce^{-\gamma t}\left\|w_{0}\right\|_{L^{1}_{k}}\qquad\text{for every $t\geq 0$}.$
This is indeed more general than our statement since one can take $w_{0}=u_{0}-F_{\varepsilon}$, but we choose to state the theorem as above for readability.
We remark that both points (i) and (ii) of the previous theorem, and all of our results concerning asymptotic behaviour, require in particular that there exists $s\in(0,1]$ such that
$\int_{\mathbb{R}^{d}}J(x)|x|^{2+s}\,\mathrm{d}x=:\rho_{2+s}<\infty.$ (3)
This quantity $\rho_{2+s}$ will be fundamental for our estimates in Section 3.
Proving an estimate like the previous one has several consequences. One of these is the following uniform-in-time result on the closeness solutions to the local and nonlocal problems, which we prove in Section 4. For readability, we state the theorem with likely non-optimal hypotheses, since optimizing it is not the main goal of this paper. It requires the following additional hypothesis on $J$:
$\int_{\mathbb{R}^{d}}J(x)x^{\alpha}\,\mathrm{d}x=0\qquad\text{for all multiindices $\alpha$ of order $|\alpha|=3$,}$ (4)
which is satisfied for example if $J$ is symmetric.
Theorem 1.3 (Nonlocal-to-local approximation).
Take $k\geq 0$ and assume $J\in L^{1}_{k+4+d}$ satisfies (1) and (4), $J\in L^{p}$ for some $p\in(1,\infty]$. Let $u$ be the mild solution to equation (NLFP) with a probability initial condition $u_{0}\in L^{1}_{k}\cap\mathcal{C}^{4}$, and such that there exist $c>0$ and $K>k+d$ with $|D^{\alpha}u_{0}(x)|\leq c\left\langle x\right\rangle^{-K}$ for all multiindices $\alpha$ with $|\alpha|=4$.
Let $v$ the (unique) integrable solution to the (local) Fokker-Planck equation (FP) with the same initial condition. Then, there exists $C>0$ such that
$\left\|u(t,\cdot)-v(t,\cdot)\right\|_{L^{1}_{k}}\leq C\varepsilon^{2}\qquad\text{for every $t\geq 0$ and $\varepsilon\in(0,1]$.}$ (5)
We emphasize that the constant $C$ is independent of $\varepsilon$.
Remark 1.4.
If we do not assume (4), then under slightly weaker assumptions ($u_{0}\in L^{1}_{k}\cap\mathcal{C}^{3}$, $J\in L^{1}_{k+3+d}$) one can show that
$\left\|u(t,\cdot)-v(t,\cdot)\right\|_{L^{1}_{k}}\leq C\varepsilon\qquad\text{for every $t\geq 0$ and $\varepsilon\in(0,1]$,}$ (6)
for some (other) $\varepsilon$-independent constant $C>0$. In order to obtain this it is enough to carry out the Taylor expansions in Section 4 to one order less.
Remark 1.5.
With the same strategy one can prove the previous statement also in Lebesgue spaces with exponential weight. Assume $J\in L^{1}_{\exp,a^{\prime}}$ and $u_{0}\in L^{1}_{\exp,a^{\prime}}\cap\mathcal{C}^{3}$. Then (6) holds also in $L^{1}_{\exp,a}$ for every $a<a^{\prime}/2$. Moreover if (4) holds and $u_{0}\in\mathcal{C}^{4}$, then equation (5) is also valid in $L^{1}_{\exp,a}$, for every $a<a^{\prime}/2$. We don’t give the proofs, since they are essentially the same as the ones in $L^{1}_{k}$ spaces.
Throughout the paper (as was used in the theorem above), we will use multi-index notation for derivatives, so if $\alpha\in\mathbb{N}^{d}$ is a multi-index we write $|\alpha|=\sum_{i}\alpha_{i}$, $x^{\alpha}=x_{1}^{\alpha_{1}}x_{2}^{\alpha_{2}}\cdots x_{d}^{\alpha_{d}}$, $\alpha!=\alpha_{1}!\alpha_{2}!\cdots\alpha_{d}!$ and $D^{\alpha}\varphi(x)=\partial_{x_{1}}^{\alpha_{1}}\partial_{x_{2}}^{\alpha_{2}}\dots\partial_{x_{d}}^{\alpha_{d}}\varphi(x)$.
We emphasize once more that, although being far from optimal, the constants we find are all explicit, since proofs are completely constructive.
We will also make a few remarks regarding the unique probability equilibrium $F_{\varepsilon}$ associated to (NLFP). Theorems 1.1 and 1.3 imply in particular that $\|F_{\varepsilon}-G\|_{L^{1}_{k}}\leq C\varepsilon^{2}$, where $G$ is the standard Gaussian probability distribution in $\mathbb{R}^{d}$. But the properties of $F_{\varepsilon}$ are quite interesting in their own right. Theorem 1.1 includes the fact that $F_{\varepsilon}$ has exponentially-decaying tails as long as $J$ has the same property. However, when $J$ is compactly supported we suspect that $F_{\varepsilon}$ has tails which decay like the Poisson distribution, that is $F_{\varepsilon}\approx e^{-a|x|\log|x|}$ for large $|x|$. We are able to show an incomplete version of this statement in Section 5.3. We suspect that in general $F_{\varepsilon}$ does not have Gaussian tails, but we haven’t been able to prove this.
The regularity of $F_{\varepsilon}$ is also interesting, as we explore in Section 5.1: independently of the regularity of $J$, $F_{\varepsilon}$ appears to be smooth everywhere except at $x=0$; and its regularity at $x=0$ depends on $\varepsilon$, not on the regularity nor tails of $J$. The equilibrium $F_{\varepsilon}$ becomes more regular as $\varepsilon\to 0$. We think a more complete study of the properties of $F_{\varepsilon}$ would be interesting for a future work.
Strategy of the proof
The central point of our results is the proof of (i) and (ii) in Theorem 1.1. The proof of Theorem 1.3 can then be obtained as a consequence with fairly standard estimates. As mentioned, the proof of Theorem 1.1 uses Harris’s theorem from probability, which requires us to show a certain positivity property of solutions of our nonlocal equation (NLFP); see Section 3.1, where we give a short introduction to this. Showing that solutions are positive for any fixed $\varepsilon>0$ is not hard; and the fact that solutions to the (local) Fokker-Planck equation (FP) are positive is well known, and a consequence of its regularization properties and the parabolic Harnack inequality (or just a consequence of its explicit solution!). However, finding a positivity result which is valid for all $\varepsilon>0$ and which does not degenerate in the limit $\varepsilon\to 0$ is not an easy task. We notice that it is linked to convergence results in the central limit theorem, since the solution $u$ to (NLFP) can be written as (see Section 2.3)
$u(t)(x)=e^{(d-\frac{1}{\varepsilon^{2}})t}\sum_{n=0}^{\infty}\Big(\frac{1}{\varepsilon^{2}}\Big)^{n}\int_{0}^{t}\int_{0}^{t_{1}}\dots\int_{0}^{t_{n-1}}J_{\varepsilon}^{t_{1},\dots,t_{n}}*u_{0}(e^{t}x)\,\mathrm{d}t_{n}\dots\,\mathrm{d}t_{1},$
where it is understood that the term for $n=0$ is just $u_{0}(e^{t}x)$, and
$J_{\varepsilon}^{t_{1},\dots,t_{n}}:=J_{\varepsilon e^{t_{1}}}*\cdots*J_{\varepsilon e^{t_{n}}}(x).$
Each of the terms represents the distribution of particles which have undergone $n$ random jumps, if we see (NLFP) as the PDE for the distribution of a jump stochastic process. Proving positivity of $u$ can then be translated to the problem of understanding the behavior of repeated convolutions of scalings of the function $J$, which can in turn be understood in terms of the central limit theorem in probability. This can be made into a rigorous proof, as we show in this paper. We think it provides an intriguing link among these different tools, and it yields a result which is otherwise non-obvious to obtain.
Previous literature and related results
We are not aware of any studies regarding the specific PDE (NLFP) except for Mischler and Tristani (2017), whose results are similar to our Theorem 1.1, but use very different techniques (our conditions on $J$ are somewhat simplified, allowing less regular and non-symmetric $J$). Our result in Theorem 1.3 regarding the nonlocal-to-local limit as $\varepsilon\to 0$ is new as far as we know. There are plenty of results on related equations, of which we give a short summary now, but we highlight that nonlocal-to-local limits which preserve the asymptotic behavior in the limit are comparatively rare, and are a current source of interest in several contexts.
The approach from nonlocal diffusion-type operators of the form JuuJ*u-u to local diffusion operators has been the subject of many previous works. The PDE $\partial_{t}u=J*u-u$ is studied in Chasseigne et al. (2006), and a review of related techniques (making strong use of explicit solutions via the Fourier transform) is given in Andreu-Vaillo et al. (2010). Generalizations of this result to similar nonlocal PDE with different integral kernels are given in Molino and Rossi (2019). The same PDE is studied in Rey and Toscani (2013), with a focus on its large-time behavior. In this case the motivation comes from both physics (since this PDE is thought to be a more appropriate diffusion model in some cases) and numerical analysis (see below for more on this). Some results on its large-time behavior can be obtained by using Fourier-type distances (Carrillo and Toscani, 2007), again by taking advantage of the fact that the PDE is explicitly solvable in Fourier variables. A different nonlocal and nonlinear version of the Fokker-Planck PDE (FP) is studied in Ignat and Rossi (2007). Its relationship to the local problem is also studied there, but the results on asymptotic behavior are not preserved when passing to the local limit. Still a different nonlocal version has been recently considered in Auricchio et al. (2023), motivated by a model in collective behavior.
We mentioned that equation (NLFP) is a good test case for ideas to study the asymptotic behavior of nonlocal PDE, and we would like to describe a few more involved situations where this is relevant. An equation close to (NLFP) was proposed in Cai et al. (2006) as a model for the probability of the concentration of a certain protein which is being expressed by a particular gene. This equation was later studied in Cañizo et al. (2019) by means of an entropy inequality developed specifically for this PDE, which highlights that a general strategy is lacking even for such simple nonlocal equations.
Another important domain where these ideas are relevant is numerical analysis. In dimension $d=1$, if we formally take $J=\frac{1}{2}\left(\delta_{-1}+\delta_{1}\right)$ then the equation $\partial_{t}u=\frac{1}{\varepsilon^{2}}(J_{\varepsilon}*u-u)$, when evaluated at a set $\{x_{i}\}_{i=0,\dots,N}$ of equally spaced nodes at distance $\varepsilon$, becomes a numerical scheme for solving the standard heat equation on $\mathbb{R}$. The proof in this paper does not apply directly to this case, since this kernel does not have the $L^{p}$ regularity we need. The question of whether a numerical method preserves the large-time behavior of its limiting equation is of practical importance, and it is a subject of recent research Ayi et al. (2022), Dujardin et al. (2020), Cancès et al. (2020) and Bessemoulin-Chatard et al. (2020). We sometimes refer to this as the property of the scheme being time asymptotic-preserving (just “asymptotic-preserving” is often used for schemes which are well-behaved under a certain scaling limit of their associated PDE, which is not our case). Numerical schemes for PDE are nonlocal (in the sense that the equation at one of the nodes depends on the value at other nodes, which are at a fixed positive distance) and the study of their asymptotic behavior often encounters similar difficulties as for nonlocal PDE. This was one of the motivations of Rey and Toscani (2013), and we have the same hope that developing different methods for nonlocal PDEs can also suggest ideas to understand the behavior of some numerical methods.
We finally mention that several methods to study the asymptotic behavior of nonlocal PDE have been developed in the context of kinetic theory. Some examples for linear PDE include semigroup tools and spectral methods (Gualdani et al., 2018), techniques for hypocoercive PDE (Villani, 2009; Dolbeault et al., 2009), applications of Harris’s theorem (Cañizo et al., 2020), and entropy techniques, which are often useful also for nonlinear PDE (Lods et al., 2008; Carrillo and Toscani, 2007; Bisi et al., 2015). Of these methods, the first one based on spectral methods is the basis of Mischler and Tristani (2017); entropy techniques are the main tool of Cañizo et al. (2019) on a very similar PDE; and as far as we know our paper is the first time Harris-based techniques are used to get the correct behavior as $\varepsilon\to 0$. It does not seem straightforward to use entropy techniques to get a similar result, since an inequality along the lines of Cañizo et al. (2019) does not respect the scaling as $\varepsilon\to 0$, but finding a way to do so is in our opinion an interesting problem.
Structure of the paper
In Section 2, we present results related to the well-posedness of the equation (NLFP) in $L_{k}^{1}$. For this purpose we define the mild solutions of (NLFP) by a standard use of Duhamel’s formula, for which we prove existence, uniqueness and continuous dependence on the initial data via the well known fixed-point strategy. We point out that the same strategy can be used to prove well-posedness in measure spaces, and interpret these results also in terms of semigroups.
In the last part of Section 2 we give two explicit representations of the solution: a classical one via Fourier transform, and the Wild sum representation which we use in a crucial way.
Section 3 contains the proof of Theorem 1.1. We give a short introduction to Harris’s Theorem, which requires two main conditions: a Lyapunov confinement condition which is fairly straightforward in our case, based on a Taylor’s expansion; and a positivity condition which is the key part of our argument, as we discussed above. Proving this positivity condition requires a suitable version of the Berry-Esseen Theorem, a form of the central limit theorem which gives a rate of convergence in the $L^{\infty}$ norm. We need a result for independent, not necessarily identically distributed, random vectors. Afterwards, the proof of positivity is straightforward once a suitable scaling is carried out.
Section 4 deals with the speed of convergence of the Nonlocal Fokker Planck equation (NLFP) to the local on (FP) at all time $t$, and gives the proof of Theorem 1.3. The rate of convergence at finite times comes from a standard consistency and stability argument in the spirit of numerical methods. The latter result, combined with Theorem 1.1, allows us to extend the convergence for all times.
In the last section 5 we gather some independent results on the regularity of the equilibrium, showing that it improves as $\varepsilon\to 0$. Moreover, we give a representation of the solution of (NLFP) in terms of the cumulative generating function, which is nothing else than the logarithm of the moment generating function. This gives in particular an explicit representation for the moments of the equilibrium. We also investigate on the tails of the solution of (NLFP) and its equilibrium $F_{\varepsilon}$. In the local case, the equilibrium is a Gaussian and the solution has Gaussian tails if the initial data has. For the nonlocal equation (NLFP) we strongly suspect this is not the case, even for very nice initial data and kernel. We are able to prove that for compactly supported $J$ the tails are at least Poisson-like.
2 Well-posedness of the initial value problem
We gather in this section some results on the well-posedness of equation (NLFP), that is,
$\partial_{t}u=\frac{1}{\varepsilon^{2}}\big(J_{\varepsilon}*u-u\big)+\operatorname{div}(xu),$
with the addition of an initial condition $u_{0}\in L^{1}_{k}$.
$u(0,x)=u_{0}(x),\qquad x\in\mathbb{R}^{d}.$
Though they are mostly standard results for linear PDE, it is useful to lay out the basic theory.
2.1 Existence and Uniqueness
Solutions to equation (NLFP) can be studied (for example) by using semigroup theory, which easily yields mild solutions $u\in\mathcal{C}([0,\infty),L_{k}^{1}(\mathbb{R}^{d}))$ whenever $u_{0}\in L^{1}_{k}$, for some $k\geq 0$. Alternatively, we give a direct simple proof of existence of mild solutions. The first observation is that the equation
$\partial_{t}u=\operatorname{div}(xu)-\frac{1}{\varepsilon^{2}}u,\qquad u(0,x)=u_{0}(x)$
has the explicit classical solution
$u(t,x)=e^{(d-1/\varepsilon^{2})t}u_{0}(e^{t}x),\qquad t\geq 0,x\in\mathbb{R}^{d},$
at least for a differentiable $u_{0}$. Hence, for any $u_{0}\in L^{1}$ we define
$T_{t}[u_{0}](x)\equiv T_{t}u_{0}(x):=e^{(d-1/\varepsilon^{2})t}u_{0}(e^{t}x)\qquad\text{for $t\geq 0$, $x\in\mathbb{R}^{d}$.}$ (7)
Following a well known strategy, we use Duhamel’s formula to define mild solutions of equation (NLFP) as follows:
Definition 2.1 (Mild solution).
Let $k\geq 0$, and take $J$ satisfying Hypothesis (1) on $J$ and $J\in L^{1}_{k}$. Take $u_{0}\in L^{1}_{k}$, and take $T^{*}\in(0,+\infty]$. A mild solution of equation (NLFP) on $0,T^{*})$ with initial condition $u_{0}$ is a function $u\in\mathcal{C}([0,T^{*}),L^{1}_{k})$ such that
$u(t)=T_{t}u_{0}+\frac{1}{\varepsilon^{2}}\int_{0}^{t}T_{t-s}\big[J_{\varepsilon}*u(s)\big]\,\mathrm{d}s\qquad\text{for all $t\in I$,}$
where the integral is understood (equivalently) in the Riemann or Bochner sense.
2.1.1 Existence of $L^{1}_{k}$ solutions
We are going to show the following result:
Theorem 2.2.
Take $k\geq 0$ and let $J$ satisfy (1) and, in addition, $J\in L_{1}^{k}$. Then, there exists a unique mild solution of equation (NLFP) with initial condition $u_{0}\in L^{1}_{k}$ (in the sense of Definition 2.1). In addition, the solution is continuously dependent on the initial data, in the following sense: for any two mild solutions $u$ and $v$ of (NLFP) with initial data $u_{0}$ and $v_{0}$
$\left\|u(t)-v(t)\right\|_{L^{1}_{k}}\leq e^{\frac{t}{\varepsilon^{2}}(c_{k}-1)}\left\|u_{0}-v_{0}\right\|_{L^{1}_{k}}$
The proof of this is fairly straightforward, and we give it after a couple of lemmas. We first notice that the convolution operator $u\mapsto J*u$ is a continuous operator under fairly mild conditions. For any measurable functions $u\colon\mathbb{R}^{d}\to\mathbb{R}$ and $V\colon\mathbb{R}^{d}\to[1,+\infty)$ we denote
$\|u\|_{V}:=\int_{\mathbb{R}^{d}}|u(x)|V(x)\,\mathrm{d}x.$
We need the following generalization of Young’s inequality in weighted spaces, whose proof we omit since it is obtained by a direct computation using the subadditivity in the assumptions:
Lemma 2.3.
Take a measurable $J\colon\mathbb{R}^{d}\to\mathbb{R}$ such that $\|J\|_{V}\|u\|_{V}<+\infty$. Then for every sub-additive function $V\colon\mathbb{R}^{d}\to[1,\infty)$ (that is, satisfying for all $x,y\in\mathbb{R}^{d}$ $V(x+y)\leq C(V(x)+V(y))$ for some $C>0$ ) it holds that
$\left\|J*u\right\|_{V}\leq 2C\left\|J\right\|_{V}\left\|u\right\|_{V}$
for all measurable $u\colon\mathbb{R}^{d}\to\mathbb{R}$ such that $\|u\|_{V}<+\infty$. In particular, for $V(x):=1+|x|^{k}$ with $k\geq 0$ we have
$\left\|J*u\right\|_{L^{1}_{k}}\leq 2^{k}\left\|J\right\|_{L^{1}_{k}}\left\|u\right\|_{L^{1}_{k}}.$ (8)
We also notice that for any $t\geq 0$, the operator $T_{t}$ is contractive in all $L^{1}_{k}$ for any $k\geq 0$:
Lemma 2.4 (Contractivity).
For all $k\geq 0$, $t>0$ and $u\in L^{1}_{k}$, the operator $T_{t}$ defined in (7) satisfies
$\|T_{t}u\|_{L^{1}_{k}}\leq e^{-t/\varepsilon^{2}}\|u\|_{L^{1}_{k}}.$
Proof.
We just write the following direct estimate:
$\|T_{t}u\|_{L^{1}_{k}}=\int(|x|^{k}+1)|T_{t}u(x)|\,\mathrm{d}x=e^{-t/\varepsilon^{2}}\int(|x|^{k}+1)|u(xe^{t})|e^{dt}\,\mathrm{d}x\\ =e^{-t/\varepsilon^{2}}\int(|e^{-t}y|^{k}+1)|u(y)|\,\mathrm{d}y<\int(|y|^{k}+1)u(y)\,\mathrm{d}y.$
Proof of Theorem 2.2.
This is done via a standard contraction argument. Take $T^{*}>0$ and define the Banach space $\mathcal{Y}:=\mathcal{C}([0,T^{*}],L^{1}_{k})$ equipped with the norm
$\left\|u\right\|_{\mathcal{Y}}:=\sup_{t\in[0,T^{*}]}\left\|u(t)\right\|_{L_{k}^{1}}.$
Let us define $\Psi:\mathcal{Y}\rightarrow\mathcal{Y}$ by
$\Psi[u](t):=T_{t}u_{0}+\frac{1}{\varepsilon^{2}}\int_{0}^{t}T_{t-s}\big[J_{\varepsilon}*u(s)\big]\,\mathrm{d}s$
for any $u\in\mathcal{Y}$, so that $u$ is a mild solution if and only if $u$ is a fixed point for $\Psi$. In the following estimate $u_{s}$ and $v_{s}$ will be used as shorthand for $u(s)$ and $v(s)$. For every $u_{t},v_{t}\in L^{1}_{k}$, use Lemmas 2.3 and 2.4 we have
$\begin{split}\|\Psi[u](t)&-\Psi[v](t)\|_{L^{1}_{k}}=\left\|\frac{1}{\varepsilon^{2}}\int_{0}^{t}T_{t-s}\big[J_{\varepsilon}*(u_{s}-v_{s})\big]\,\mathrm{d}s\right\|_{L^{1}_{k}}\\ &\leq\frac{1}{\varepsilon^{2}}\int_{0}^{t}\left\|T_{t-s}\big[\big|J_{\varepsilon}*(u_{s}-v_{s})\big|\big]\right\|_{L^{1}_{k}}\,\mathrm{d}s\\ &\leq\frac{1}{\varepsilon^{2}}\int_{0}^{t}e^{-\frac{t-s}{\varepsilon^{2}}}\left\|J_{\varepsilon}*(u_{s}-v_{s})\right\|_{L^{1}_{k}}\,\mathrm{d}s\\ &\leq\frac{1}{\varepsilon^{2}}c_{k}\int_{0}^{t}\|u_{s}-v_{s}\|_{L_{k}^{1}}\,\mathrm{d}s\\ &\leq\frac{1}{\varepsilon^{2}}c_{k}t\,\,\|u-v\|_{\mathcal{Y}},\end{split}$
where $c_{k}=2^{k}\left\|J\right\|_{L^{1}_{k}}\geq 2^{k}\left\|J_{\varepsilon}\right\|_{L^{1}_{k}}$ from Lemma 2.3. Iterating the process, we obtain
$\|\Psi^{n}[u](t)-\Psi^{n}[v](t)\|_{L^{1}_{k}}\leq\Big(\frac{c_{k}t}{\varepsilon^{2}}\Big)^{n}\frac{1}{n!}\|u-v\|_{\mathcal{Y}}\leq\Big(\frac{c_{k}T^{*}}{\varepsilon^{2}}\Big)^{n}\frac{1}{n!}\|u-v\|_{\mathcal{Y}}$
and then, taking the supremum, leads to
$\begin{split}\|\Psi^{n}[u](t)-\Psi^{n}[v](t)\|_{\mathcal{Y}}\leq\Big(\frac{c_{k}T^{*}}{\varepsilon^{2}}\Big)^{n}\frac{1}{n!}\|u-v\|_{\mathcal{Y}}\end{split}$
For every $T^{*}$, there exists $n\in\mathbb{N}$ such that $\big(\frac{c_{k}T^{*}}{\varepsilon^{2}}\big)^{n}\frac{1}{n!}<1$. There is a specific version of the contraction mapping theorem (see for example Miklavcic (1998, Thm. 1.1.3)) which now implies $\Psi$ has a fixed point which is a solution of (NLFP) in the sense of Definition 2.1. (Alternatively, one can consider the sequence $(\Psi^{n}(u_{0}))_{n\geq 0}$, where $u_{0}\in\mathcal{Y}$ is the function constantly equal to $u_{0}$ on $0,T^{*}]$; our last inequality easily shows that it is a Cauchy sequence, and hence must have a limit. This can be done for any $T^{*}>0$, easily obtaining the result on $0,+\infty)$.)
To prove continuity with respect to initial data we proceed as before: let $v$ be another mild solution for some other initial data $v_{0}$, then for every $t$
$\begin{split}&\left\|u(t)-v(t)\right\|_{L^{1}_{k}}\leq\left\|T_{t}(u_{0}-v_{0})\right\|_{L^{1}_{k}}+\frac{1}{\varepsilon^{2}}\int_{0}^{t}\left\|T_{t-s}[J_{\varepsilon}*(u_{s}-v_{s})]\right\|_{L^{1}_{k}}\,\mathrm{d}s\\ &\leq e^{-t/\varepsilon^{2}}\left\|u_{0}-v_{0}\right\|_{L^{1}_{k}}+\frac{c_{k}}{\varepsilon^{2}}\int_{0}^{t}e^{-\frac{t-s}{\varepsilon^{2}}}\left\|u(s)-v(s)\right\|_{L^{1}_{k}}\,\mathrm{d}s.\end{split}$
By the standard Gronwall lemma applied to $e^{\frac{t}{\varepsilon^{2}}}\left\|u(t)-v(t)\right\|_{L^{1}_{k}}$ we obtain the result. ∎
2.1.2 Existence of measure solutions
We sketch similar arguments that give existence of solutions in a space of measures, since Harris’s theorem we use later is usually stated for semigroups in a space of measures (or Markov semigroups in the context of probability). This is really not strictly necessary, since there are versions of the Harris theorem which hold similarly in weighted $L^{1}$ spaces (see Cañizo and Mischler (2023)), but since the proof is very similar we give the main steps.
Let us denote with $\mathcal{M}(\mathbb{R}^{d})$ the space of signed finite Radon measures over $\mathbb{R}^{d}$ and by $\left\|\cdot\right\|_{TV}$ and $\left\|\cdot\right\|_{\mathcal{M}_{k}}$ the total variation norm and the weighted total variation, respectively. Let us recall the definition of Bounded Lipschitz norm, i.e.
$\left\|\mu\right\|_{BL}:=\sup_{f\in\mathcal{F}}\left|\int_{\mathbb{R}^{d}}f\,\mathrm{d}\mu\right|$
where $\mathcal{F}:=\{f\in W^{1,\infty}(\mathbb{R}^{d}):\left\|f\right\|_{1,\infty}\leq 1\}$, with $\left\|f\right\|_{1,\infty}:=\|f\|_{\infty}+\operatorname{Lip}(f)$ and $\operatorname{Lip}(f)$ denoting the Lipschitz constant of $f$. In the literature, the bounded Lipschitz norm is sometimes defined as the equivalent norm obtained by choosing in the previous definition $\mathcal{F}:=\{f:\mathbb{R}^{d}\to\mathbb{R}:\left\|f\right\|_{\infty}\leq 1,\sup_{x\neq y}\frac{\left|f(x)-f(y)\right|}{|x-y|}\leq 1\}$.
We define the ball $\mathcal{X}_{R}:=\{\mu\in\mathcal{M}:\left\|\mu\right\|_{\mathcal{M}_{k}}\leq R\}$ and the set $\mathcal{Z}=\mathcal{C}([0,T^{*}],\mathcal{X}_{R})$ of functions which are continuous from $0,T^{*}]$ to $\mathcal{X}_{R}$ in the bounded Lipschitz norm. $\mathcal{Z}$ is a complete metric space with the metric induced by the norm
$\left\|\mu\right\|_{\mathcal{Z}}=\sup_{t\in[0,T^{*}]}\left\|\mu(t)\right\|_{BL}.$
We stress that the same claim in the space $\mathcal{C}([0,\infty),\mathcal{M}_{k})$ does not hold since in $(W^{1,\infty})^{*}$ there are elements that are not measures which can be approximated by Cauchy sequences in $\mathcal{M}_{k}$.
Following the path of the previous proof, we can prove existence and uniqueness of the mild solution of equation (NLFP) in $\mathcal{Z}$ choosing a suitable $R$ such that $\mu_{0}\in\mathcal{X}_{R}$. One can prove a version of inequality (8) in weighted total variation norms:
$\left\|J*\mu\right\|_{\mathcal{M}_{k}}\leq c_{k}\left\|\mu\right\|_{\mathcal{M}_{k}}$
Moreover, the contractivity Lemma 2.4 still holds in the sense of measures in weighted total variation norm. The functional $\Psi$, defined as
$\Psi[\mu](t)=T_{t}\mu_{0}+\frac{1}{\varepsilon^{2}}\int_{0}^{t}T_{t-s}[J_{\varepsilon}*\mu_{s}]\,\mathrm{d}s$
is such that if $t\to\mu(t)\in\mathcal{Z},$ then so is $t\to\Psi[\mu](t)$ for small enough $T^{*}$. We briefly show a proof of this. Continuity comes straightforwardly from the fact that $T_{t}$ is a continuous semigroup (in the bounded Lipschitz norm) and all the other operations are continuity-preserving.
We have left to show that $\Psi[\mu](t)\in\mathcal{X}_{R}$ for every $t\in[0,T^{*}]$ when $\mu(t)\in\mathcal{X}_{R}$. Set $R=2\left\|\mu_{0}\right\|_{\mathcal{M}_{k}}$
$\begin{split}&\left\|\Psi[\mu](t)\right\|_{\mathcal{M}_{k}}\leq\left\|T_{t}\mu_{0}\right\|_{\mathcal{M}_{k}}+\frac{1}{\varepsilon^{2}}\int_{0}^{t}\left\|T_{t-s}[J_{\varepsilon}*\mu_{s}]\,\mathrm{d}s\right\|_{\mathcal{M}_{k}}\\ &\leq\left\|\mu_{0}\right\|_{\mathcal{M}_{k}}+\frac{1}{\varepsilon^{2}}\int_{0}^{t}C_{k}\left\|\mu_{s}\right\|_{\mathcal{M}_{k}}\,\mathrm{d}s\leq\left\|\mu_{0}\right\|_{\mathcal{M}_{k}}+\frac{1}{\varepsilon^{2}}CT^{*}\left\|\mu\right\|_{\mathcal{\mathcal{M}}_{k}}\leq R\end{split}$
for $T^{*}\leq\frac{\varepsilon^{2}}{2C_{k}}$. We then iterate the process for another time $T^{*}_{1}$ until we extend the existence and uniqueness globally.
With the same approach, one can prove the previous results in the space $L^{1}_{\exp,a}$. Since the procedure is the same mutatis mutandis, we have chosen to omit it.
2.2 Semigroup results
Our result on the well-posedness of the nonlocal Fokker-Planck equation (NLFP) can also be phrased in terms of semigroup theory. For an introduction and further details on the results we will be using below, we refer the reader to the books Engel and Nagel (2001) or Pazy (1983). We first point out that for any $k\geq 0$, the operators $(T_{t})_{t\geq 0}$ given in (7) define a strongly continuous semigroup on the Banach space $L^{1}_{k}$; it is easy to check this directly. The generator of this semigroup is the operator $B_{\varepsilon}$ given by
$B_{\varepsilon}u:=\operatorname{div}(xu)-\frac{1}{\varepsilon^{2}}u$
with domain
$\mathcal{D}(B_{\varepsilon}):=\{u\in L^{1}_{k}\mid\operatorname{div}(xu)\in L^{1}_{k}\},$
which is a closed linear operator. Then, equation (NLFP) can be written as
$\frac{\mathrm{d}}{\mathrm{d}t}u=L_{\varepsilon}u=A_{\varepsilon}u+B_{\varepsilon}u,$ (9)
where $L_{\varepsilon}$ and $A_{\varepsilon}$ denote the operators
$A_{\varepsilon}u:=\frac{1}{\varepsilon^{2}}J_{\varepsilon}*u,$
$L_{\varepsilon}u:=A_{\varepsilon}u+B_{\varepsilon}u=\frac{1}{\varepsilon^{2}}\left(J_{\varepsilon}*u-u\right)+\operatorname{div}(xu).$
Due to Lemma 2.3 we know that $A_{\varepsilon}\colon L^{1}_{k}\to L^{1}_{k}$ is a bounded linear operator. Since we know $B_{\varepsilon}$ is a generator in $L^{1}_{k}$, standard results for bounded perturbations in semigroup theory (see for example Engel and Nagel (2001, chapter III.1)) ensure then that the operator $L_{\varepsilon}=A_{\varepsilon}+B_{\varepsilon}$ is also a generator with domain $\mathcal{D}(L_{\varepsilon})=\mathcal{D}(B_{\varepsilon})$. This is equivalent to the well-posedness of equation (9) in the sense of semigroups (Engel and Nagel, 2001, Chapter II.6). Solutions to (9) in the sense of semigroups must also be mild solutions in the sense of Definition 2.1 (Engel and Nagel, 2001, Section III.1.3, Corollary 1.4), so they coincide with the solutions given in Theorem 2.2.
2.3 Representation of the solution
For this kind of problems, we can express the solution via the classic Fourier transform or with the so-called Wild sum. We use the latter as an important tool in our later proofs, emphasizing that our main results do not seem easy to obtain by using the explicit representation in the terms of the Fourier transform. More importantly, we would like to explore strategies which would be valid even in cases in which Fourier transform methods are not available.
For the sake of completeness and to obtain regularity results in section 5.1, we will also briefly comment on the Fourier solution. From (1) we deduce that
$\hat{J}(\xi)=1-\xi^{2}+o(|\xi|^{2})\qquad\text{as $\xi\to 0$.}$
Then if we take $u_{0}$ as an initial probability distribution in $L^{1}$, we can prove that there exists at most one solution of (NLFP) in $\mathcal{C}([0,\infty),L^{1})$. If it exists, it must satisfy
$\hat{u}(t,\xi)=\hat{u}_{0}(e^{-t}\xi)e^{\frac{1}{\varepsilon^{2}}\int_{0}^{t}\zeta_{\varepsilon}(e^{-s}\xi)\,\mathrm{d}s}$ (10)
with $\zeta_{\varepsilon}(\xi)=\hat{J}_{\varepsilon}(\xi)-1$. Moreover, there exists at most one equilibrium in $L^{1}$ of (NLFP), and it must satisfy
$\hat{F}_{\varepsilon}(\xi)=e^{\frac{1}{\varepsilon^{2}}\int_{0}^{\infty}\zeta_{\varepsilon}(e^{-s}\xi)\,\mathrm{d}s}$ (11)
Indeed, passing to Fourier variables, equation (NLFP) becomes
$\partial_{t}\hat{u}=\frac{1}{\varepsilon^{2}}\big(\hat{J}_{\varepsilon}\hat{u}-\hat{u}\big)-\xi\cdot\nabla_{\xi}\hat{u}=\frac{1}{\varepsilon^{2}}\zeta_{\varepsilon}(\xi)\hat{u}-\xi\cdot\nabla_{\xi}\hat{u}$
which has equation (10) as its unique solution. We also notice that the transform of the equilibrium must satisfy
$\hat{F_{\varepsilon}}(\xi)=\hat{u}(r,\xi)=\hat{F}_{\varepsilon}(e^{-r}\xi)e^{\frac{1}{\varepsilon^{2}}\int_{0}^{r}\zeta_{\varepsilon}(e^{-s}\xi)\,\mathrm{d}s}\quad\text{ for every $r\geq 0$}$
with the initial condition $\hat{F}_{\varepsilon}(0)=1$, since $\left\|F_{\varepsilon}\right\|_{L^{1}}=1$. In particular
$\hat{F}_{\varepsilon}(\xi)=\lim_{r\to\infty}\hat{F}_{\varepsilon}(e^{-r}\xi)e^{\frac{1}{\varepsilon^{2}}\int_{0}^{r}\zeta_{\varepsilon}(e^{-s}\xi)\,\mathrm{d}s}=e^{\frac{1}{\varepsilon^{2}}\int_{0}^{\infty}\zeta_{\varepsilon}(e^{-s}\xi)\,\mathrm{d}s}.$
Remark 2.5.
With a more careful study of these arguments or other classical ones (like Tychonoff’s Theorem) one could prove existence of a solution and of the equilibrium. We don’t want to insist on these details since in the next section we will give a different proof for the existence and uniqueness of the equilibrium.
Wild Sum
When dealing with linear convolution-type operators, Wild Sum—or Dyson-Phillips ’s Series (Engel and Nagel, 2001, p. 163 )—give a particularly simple way to express the solution, based on a “Picard iteration” for PDEs.
We first present the following lemma that, for a fixed $t$, allows us to exchange the order of the operator $T_{t}$ with the convolution operator and vice versa.
Lemma 2.6.
Let $(T_{t})$ be the semigroup associated to the solution of equation (7). Then
$J_{\varepsilon}*(T_{t}f)=T_{t}[J_{\varepsilon e^{t}}*f]$
Proof.
For every $x$
$\big(J_{\varepsilon}*T_{t}f\big)(x)=\int\varepsilon^{-d}J\Big(\frac{x-y}{\varepsilon}\Big)(T_{t}f)(y)\,\mathrm{d}y=e^{(d-\frac{1}{\varepsilon^{2}})t}\int\varepsilon^{-d}J\Big(\frac{x-y}{\varepsilon}\Big)f(e^{t}y)\,\mathrm{d}y\\ =e^{-\frac{t}{\varepsilon^{2}}}\int\varepsilon^{-d}J\Big(\frac{e^{-t}}{\varepsilon}(e^{t}x-z)\Big)f(z)\,\mathrm{d}z\\ =\int e^{(d-\frac{1}{\varepsilon^{2}})t}\big(e^{-dt}\varepsilon^{-d}\big)J\Big(\frac{e^{t}x-z}{e^{t}\varepsilon}\Big)f(z)\,\mathrm{d}z=e^{(d-\frac{1}{\varepsilon^{2}})t}\big(J_{\varepsilon e^{t}}*f\big)(e^{t}x)$ (12)
where we have used the change of variable $z=e^{t}y$. Hence,
$J_{\varepsilon}*T_{t}f=T_{t}[J_{\varepsilon e^{t}}*f].\qed$
Theorem 2.7.
Let $u$ be the unique solution to equation (NLFP) with initial condition $u_{0}\in L^{1}$. Then, in the hypotheses of Theorem 2.2, we can write the solution of the equation via Wild sum, namely for every $t\geq 0$
$u(t)(x)=e^{(d-\frac{1}{\varepsilon^{2}})t}\Big[u_{0}(e^{t}x)+\sum_{n=1}^{\infty}\Big(\frac{1}{\varepsilon^{2}}\Big)^{n}\int_{0}^{t}\int_{0}^{t_{1}}\dots\int_{0}^{t_{n-1}}J_{\varepsilon}^{t_{1},\dots,t_{n}}*u_{0}(e^{t}x)\,\mathrm{d}t_{n}\dots\,\mathrm{d}t_{1}.\Big]$
where the convergence in the RHS is understood to be in $L^{1}$ for every $t$, and
$J_{\varepsilon}^{t_{1},\dots,t_{n}}:=J_{\varepsilon e^{t_{1}}}*\cdots*J_{\varepsilon e^{t_{n}}}.$
for $0\leq t_{1}\leq\dots\leq t_{n}\leq t$.
Proof.
Define
$\begin{split}h(t):=T_{-t}u(t)&=T_{-t}\Big[T_{t}u_{0}+\int_{0}^{t}T_{t-s}[\frac{1}{\varepsilon^{2}}J_{\varepsilon}*u(s)]\,\mathrm{d}s\Big]\\ &=u_{0}+\frac{1}{\varepsilon^{2}}\int_{0}^{t}T_{-s}[J_{\varepsilon}*u(s)]\end{split}$
Using Lemma 2.6
$T_{-s}[J_{\varepsilon}*u_{s}]=T_{-s}[J_{\varepsilon}*T_{s}(T_{-s}u)]=T_{-s}[T_{s}J_{\varepsilon e^{s}}*h(s)]=J_{\varepsilon e^{s}}*h(s)$
This means that $u$ is a mild solution for Equation (NLFP) iff
$h(t)=u_{0}+\frac{1}{\varepsilon^{2}}\int_{0}^{t}J_{\varepsilon e^{s}}*T_{-s}u(s)\,\mathrm{d}s=u_{0}+\frac{1}{\varepsilon^{2}}\int_{0}^{t}J_{\varepsilon e^{s}}*h(s)\,\mathrm{d}s,$
that is if $h$ is a fixed point for the functional $\Psi$
$\Psi(g):=u_{0}+\frac{1}{\varepsilon^{2}}\int_{0}^{t}J_{\varepsilon e^{s}}*g_{s}\,\mathrm{d}s$
To find an expression for this fixed point we perform a Picard Iteration defining $h_{0}=u_{0}$ and
$h_{n+1}=\Psi(h_{n})$ (13)
i.e.
$h_{n}(t)=h_{n-1}(t)+\frac{1}{\varepsilon^{2n}}\int_{0}^{t}\int_{0}^{t_{1}}\dots\int_{0}^{t_{n-1}}J_{\varepsilon}^{t_{1},\dots,t_{n}}*u_{0}\,\mathrm{d}t_{n}\dots\,\mathrm{d}t_{1}.$
Clearly the sequence $(h_{n})_{n}$ is pointwise monotone increasing. Since the functions $h_{n}$ are all nonnegative and bounded above by $h$, the monotone convergence theorem shows that $h_{n}$ converges in $L^{1}$ to some function $\tilde{h}=\tilde{h}(t)\geq 0$ when $n\rightarrow\infty$, namely
$\tilde{h}(t)=u_{0}+\sum_{n=1}^{\infty}\Big(\frac{1}{\varepsilon^{2}}\Big)^{n}\int_{0}^{t}\int_{0}^{t_{1}}\dots\int_{0}^{t_{n-1}}J_{\varepsilon}^{t_{1},\dots,t_{n}}*u_{0}\,\mathrm{d}t_{n}\dots\,\mathrm{d}t_{1}.$
Passing to the limit in (13), one obtains that $\tilde{h}$ is a fixed point for the functional $\Psi$, so in fact $\tilde{h}=h$. Therefore,
$u(t)(x)=\big(T_{t}h\big)(x)=e^{(d-1/\varepsilon^{2})t}u_{0}(e^{t}x)+$
$+e^{(d-1/\varepsilon^{2})t}\sum_{n=1}^{\infty}\varepsilon^{-2n}\int_{0}^{t}\int_{0}^{t_{1}}\dots\int_{0}^{t_{n-1}}\Big(J_{\varepsilon}^{t_{1},\dots,t_{n}}*u_{0}\Big)(e^{t}x)\,\mathrm{d}t_{n}\dots\,\mathrm{d}t_{1}.\qed$
Remark 2.8.
From the previous it is immediate to see that if $u_{0}$ and $J$ are radially symmetric functions then also the solution is radially symmetric. One can also say this of any property which is preserved by convolution with $J$. Analogously, the equilibrium will preserve the same properties.
3 Asymptotic behavior
In this section we study the asymptotic behavior of equation (NLFP). For this we use Harris’s theorem from probability, establishing exponential convergence to equilibrium of solutions in $\mathcal{M}_{k}$ for $k>0$. Our crucial advantage of our approach is that it yields a uniform bound on the convergence rate to equilibrium for all $\varepsilon\in(0,1]$.
3.1 Harris’s Theorem
Let us first give a short overview of Harris’s theorem, in order to make this paper as self-contained as possible. We refer again to Hairer and Mattingly (2011); Cañizo and Mischler (2023) and the book by Meyn and Tweedie (2010) for a more complete exposition.
Let us denote by $\mathcal{M}\equiv\mathcal{M}(\Omega)$, the space of finite signed measures over a measurable space $\Omega$, with $\left\|\mu\right\|=\left\|\mu\right\|_{TV}=\int|\mu|$, the total variation norm. $\mathcal{P}$ will denote the set of probability measures in $\mathcal{M}$. If $V:\Omega\rightarrow[1,\infty)$ is a measurable function (which we call a weight function), we denote by $\mathcal{M}_{V}$ the subspace of measures $\mu\in\mathcal{M}$ such that the norm $\left\|\mu\right\|_{V}=\int V|\mu|<\infty$. We will use $\mathcal{P}_{V}$ to denote $\mathcal{M}_{V}\cap\mathcal{P}$.
A linear operator $S:\mathcal{M}\to\mathcal{M}$ is stochastic if it leaves $\mathcal{P}$ invariant. A stochastic semigroup $(S_{t})_{t\geq 0}$ is a family of stochastic operators $S_{t}$ parameterised by $t$ such that $S_{0}={\textrm{Id}}\,$ and $S_{t}\circ S_{s}=S_{t+s}$ for all $t,s\geq 0$.
We say that $S$ is a stochastic operator on $\mathcal{M}_{V}$ if it is a stochastic operator on $\mathcal{M}$ and its restriction on $\mathcal{M}_{V}$ is bounded in the $\left\|\cdot\right\|_{V}$ norm. Similarly, $S_{t}$ is a stochastic semigroup on $\mathcal{M}_{V}$ if it is a stochastic semigroup on $\mathcal{M}$ and each $S_{t}$ is a stochastic operator on $\mathcal{M}_{V}$.
A basic tool in the study of the asymptotic behavior of Markov Processes is Doeblin’s Theorem that under a (strong) positivity hypothesis gives existence of a unique stationary state, which is additionally exponentially stable. A well known generalization of it is Harris’s Theorem where Doeblin’s condition is replaced by a weaker local positivity condition and a Lyapunov confinement, which we state next.
Hypothesis 1 (Semigroup Lyapunov condition).
Let $V\colon\Omega\to[1,\infty)$ be a measurable (“weight”) function and let $(S_{t})_{t\geq 0}$ be a stochastic semigroup on $\mathcal{M}_{V}$. We say that $(S_{t})_{t\geq 0}$ satisfies a semigroup Lyapunov condition if there exist constants $\lambda$, $C>0$ such that
$\left\|S_{t}\mu\right\|_{V}\leq e^{-\lambda t}\left\|\mu\right\|_{V}+\frac{C}{\lambda}(1-e^{-\lambda t})\left\|\mu\right\|.$ (SLC)
If $S_{t}$ is a semigroup with generator $L$, a natural and better known version of this condition is to impose
$L^{*}V\leq C-\lambda V$ (14)
where $L^{*}$ is the (formal) dual operator of $L$ acting on “nice enough” functions (see for example Hairer (2010), where inequality (14) must be understood in a submartingale sense). We show condition (14), in a suitable sense, for the generator associated to equation (NLFP) in Theorems 3.2 and 3.4. From this we will deduce Hypothesis 1.
The second condition needed is the Harris condition:
Hypothesis 2 (Harris condition).
A stochastic operator $S$ on $\mathcal{M}$ satisfies a Harris condition on a set $\mathcal{S}\subseteq\Omega$ if there exist $0<\alpha<1$ and $\rho\in\mathcal{P}$ such that
$S\mu\geq\alpha\rho\int_{\mathcal{S}}\mu\quad\text{ for every }0\leq\mu\in\mathcal{M}.$ (HC)
We will use the following version of Harris’s Theorem:
Theorem 3.1 (Harris’s theorem for semigroups.).
Let $V:\Omega\to[1,\infty)$ be a measurable function and let $(S_{t})_{t\geq 0}$ be a stochastic semigroup on $\mathcal{M}_{V}$. Assume that:
  1. 1.
    the semigroup $(S_{t})_{t\geq 0}$ satisfies the semigroup Lyapunov condition (SLC), and
  2. 2.
    for some $T^{*}>0$ and some $R>\frac{2C}{\lambda}$, the operator $S_{T^{*}}$ satisfies the Harris condition (HC) on the set $\mathcal{S}:=\{x:V(x)\leq R\}$.
Then the semigroup has an invariant probability $\mu^{*}\in\mathcal{P}_{V}$ which is unique in $\mathcal{P}_{V}$, and there exist $C,\lambda>0$ such that
[…]
and in particular
$\left\|S_{t}\mu-\mu^{*}\right\|_{V}\leq Ce^{-\lambda t}\left\|\mu-\mu^{*}\right\|_{V}\quad\text{ for all }\mu\in\mathcal{P}_{V}\text{ and }t\geq 0.$
Short and elementary proofs of Harris’s theorem are available (Hairer and Mattingly, 2011; Cañizo and Mischler, 2023). The theorem just makes rigorous the fact that a uniform positivity condition plus a confinement condition imply exponential convergence to equilibrium. The central part of any arguments using Harris’s theorem must then be contained in the proof of these conditions, and usually (like in our case) in the proof of the positivity condition.
Theorem 1.1 automatically follows once we prove the two conditions of Harris’s Theorem with weight $V(x)=\left\langle x\right\rangle^{k}$. We notice that, since solutions to (NLFP) with $L^{1}_{k}$ initial data remain in $L^{1}_{k}$ for all times (and the $\mathcal{M}_{k}$ norm in $L^{1}_{k}$ is just equal to the $L^{1}_{k}$ norm), Theorem 1.1 in $L^{1}_{k}$ really follows once we are able to apply Theorem 3.1 in $\mathcal{M}_{k}$. A similar consideration holds with exponential weights. Hence, the next two sections are devoted to proving the hypotheses required in Theorem 3.1 for the semigroup generated by equation (NLFP).
3.2 Lyapunov condition
We define $L_{\varepsilon}^{*}$, the formal dual operator of $L_{\varepsilon}$, acting on functions $\varphi\in\mathcal{C}^{2}$ (with appropriate growth as $|x|\to+\infty$) as
$L_{\varepsilon}^{*}\varphi:=A^{*}_{\varepsilon}\varphi+B_{\varepsilon}^{*}\varphi=\frac{1}{\varepsilon^{2}}\int_{\mathbb{R}^{d}}J_{\varepsilon}(y)(\varphi(x+y)-\varphi(x))\,\mathrm{d}y-x\nabla\varphi.$
This “appropriate growth” is just to ensure the convolution $J_{\varepsilon}*\varphi$ is properly defined. Let $(S_{t})_{t\geq 0}$ be the semigroup associated to equation (NLFP).
Theorem 3.2.
Let $J$ satisfy (1) and assume $J\in L^{1}_{k^{\prime}}$ for some $k^{\prime}\geq 2$. Then for every $0<k\leq k^{\prime}$ there exist constants $C,\lambda>0$ depending only on $k$ and $J$ such that
$L_{\varepsilon}^{*}(\left\langle x\right\rangle^{k})\leq C-\lambda\left\langle x\right\rangle^{k}$
for any $0<\varepsilon\leq 1$.
Proof.
To lighten the notation in the following, we will use $C_{1},C_{2}$ as general constants, whose value may change from line to line. Recall from (1), that for all $i=1,\dots,d$
$\int J(z)z_{i}\,\mathrm{d}z=0.$
Therefore, for every $\varphi\in\mathcal{C}^{\infty}$, expanding around $x$,
$\begin{split}\frac{1}{\varepsilon^{2}}\int_{\mathbb{R}^{d}}J_{\varepsilon}(y)(\varphi(x+y)-\varphi(x))\,\mathrm{d}y&=\frac{1}{\varepsilon^{2}}\int_{\mathbb{R}^{d}}J(z)(\varphi(x+\varepsilon z)-\varphi(x))\,\mathrm{d}z\\ &=\frac{1}{\varepsilon^{2}}\int_{\mathbb{R}^{d}}J(z)\Big(\varphi(x)+\frac{1}{2}\sum_{|\alpha|=2}R_{\alpha}(\xi)\varepsilon^{2}z^{\alpha}-\varphi(x)\Big)\,\mathrm{d}z\\ &=\frac{1}{2}\int_{\mathbb{R}^{d}}J(z)\sum_{i=1}^{d}\partial_{i}^{2}\varphi(\xi)z_{i}^{2}\,\mathrm{d}z\end{split}$
where $R_{\alpha}(\xi)$ is the Lagrange reminder for some $\xi(x,z)\in$ $x,x+\varepsilon z]$, the segment that joins $x$ and $x+\varepsilon z$.
Fix now $\varphi(x)=\left\langle x\right\rangle^{k}$. With a simple computation
$\partial_{i}^{2}\left\langle x\right\rangle^{k}=k\big(\left\langle x\right\rangle^{k-2}+(k-2)x_{i}^{2}\left\langle x\right\rangle^{k-4}\big).$ (15)
For $k\in(0,2)$ we have $\partial_{i}^{2}\varphi(\xi)\leq k$, for every $i,\xi$ . Hence
$\frac{1}{\varepsilon^{2}}\int_{\mathbb{R}^{d}}J_{\varepsilon}(y)(\left\langle x+y\right\rangle^{k}-\left\langle x\right\rangle^{k})\,\mathrm{d}y\leq\frac{k}{2}\int_{\mathbb{R}^{d}}J(z)|z|^{2}\,\mathrm{d}z=k.$
For $k\geq 2$, since $\left\langle\cdot\right\rangle^{j}$ is a radially nondecreasing subadditive function for every $j\geq 0$, we can bound (15) by
$\begin{split}\partial_{i}^{2}\varphi(\xi)&\leq k(k-1)\langle\xi\rangle^{k-2}\leq k(k-1)\left\langle|x|+\varepsilon|z|\right\rangle_{k-2}\leq C_{1}(\left\langle x\right\rangle^{k-2}+\left\langle z\right\rangle^{k-2}).\end{split}$
It follows that
$\begin{split}\frac{1}{\varepsilon^{2}}&\int_{\mathbb{R}^{d}}J_{\varepsilon}(y)(\left\langle x+y\right\rangle^{k}-\left\langle x\right\rangle^{k})\,\mathrm{d}y\leq\frac{1}{2}\int_{\mathbb{R}^{d}}J(z)\sum_{i=1}^{d}z_{i}^{2}C_{1}(\langle x\rangle^{k-2}+\langle z\rangle^{k-2})\,\mathrm{d}z\\ &\leq C_{1}\langle x\rangle^{k-2}\frac{1}{2}\int_{\mathbb{R}^{d}}J(z)|z|^{2}\,\mathrm{d}z+C_{1}\int_{\mathbb{R}^{d}}J(z)\langle z\rangle^{k}\,\mathrm{d}z\leq C_{2}\langle x\rangle^{k-2}.\end{split}$
Consequently, for every $k\geq 0$
$\begin{split}L^{*}_{\varepsilon}&\left\langle x\right\rangle^{k}=\frac{1}{\varepsilon^{2}}\int_{\mathbb{R}^{d}}J_{\varepsilon}(y)(\left\langle x+y\right\rangle^{k}-\left\langle x\right\rangle^{k})\,\mathrm{d}y-x\nabla\varphi\leq C_{2}\left\langle x\right\rangle^{k-2}-k|x|^{2}\left\langle x\right\rangle^{k-2}\\ &=(C_{2}+k)\left\langle x\right\rangle^{k-2}-k(1+|x|^{2})\left\langle x\right\rangle^{k-2}\leq C-\lambda\left\langle x\right\rangle^{k}\end{split}$ (16)
for some $C,\lambda>0$. ∎
Remark 3.3.
We point out that as $k$ approaches 0, the right hand side of (16) goes to 0 as well, since we don’t have Lyapunov confinement in $L^{1}$ without weight.
With almost the same procedure, we can prove a Lyapunov condition in $L^{1}_{\exp,a}$.
Theorem 3.4.
Assume $J\in L^{1}_{\exp,a^{\prime}}$ for some $a^{\prime}>0$. Then for every $0<a<a^{\prime}$ there exist $C,\lambda>0$ (only depending on $a$, $d$ and $J$), such that
$L_{\varepsilon}^{*}(e^{a\left\langle x\right\rangle})\leq C-\lambda e^{a\left\langle x\right\rangle}.$
Proof.
To lighten the notation in the following, we will use $C_{1},C_{2}$ as general constants, whose value may change from line to line. As before, for all $i=1,\dots,d$
$\int_{\mathbb{R}^{d}}J(z)z_{i}\,\mathrm{d}z=0.$
As before, for every $\varphi\in\mathcal{C}^{\infty}$, we Taylor expand around $x$, obtaining
$\begin{split}\frac{1}{\varepsilon^{2}}\int_{\mathbb{R}^{d}}J_{\varepsilon}(y)(\varphi(x+y)-\varphi(x))\,\mathrm{d}y=\frac{1}{2\varepsilon^{2}}\int_{\mathbb{R}^{d}}J(z)\sum_{i=1}^{d}\partial_{i}^{2}\varphi(\xi)\varepsilon^{2}z_{i}^{2}\,\mathrm{d}z\end{split}$
where the Lagrange reminder is evaluated at some $\xi(x,z)\in$ $x,x+\varepsilon z]$.
Set now $\varphi=e^{a\left\langle x\right\rangle}$. One can see that there exists $C_{1}>0$ a dimension dependent constant, such that
$\partial_{i}^{2}\varphi(\xi)=\frac{ae^{a\left\langle\xi\right\rangle}}{\left\langle\xi\right\rangle^{2}}\Big(a\xi_{i}^{2}+\left\langle\xi\right\rangle-\frac{\xi_{i}^{2}}{\left\langle\xi\right\rangle}\Big)\leq C_{1}e^{a\left\langle\xi\right\rangle}\leq C_{1}e^{a(\left\langle x\right\rangle+\varepsilon\left\langle z\right\rangle)}$
since $\varphi$ is a symmetric, positive and increasing function in every component and $\left\langle\xi\right\rangle\leq\langle|x|+\varepsilon|y|\rangle\leq\left\langle x\right\rangle+\varepsilon\langle y\rangle$. Thus
$\begin{split}\frac{1}{\varepsilon^{2}}\int_{\mathbb{R}^{d}}J_{\varepsilon}(y)(\varphi(x+y)-\varphi(x))\,\mathrm{d}y&\leq\frac{1}{2}\int_{\mathbb{R}^{d}}J(z)\sum_{i=1}^{d}C_{1}e^{\langle|x|+\varepsilon|z|\rangle}z_{i}^{2}\,\mathrm{d}z\\ &\leq\frac{C_{1}}{2}e^{a\left\langle x\right\rangle}\int_{\mathbb{R}^{d}}J(z)e^{a\varepsilon\left\langle z\right\rangle}|z|^{2}\,\mathrm{d}z\leq C_{2}e^{a\left\langle x\right\rangle}\end{split}$
since
$\int_{\mathbb{R}^{d}}J(z)|z|^{2}e^{a\varepsilon\left\langle z\right\rangle}\,\mathrm{d}z\leq C_{3}\int_{\mathbb{R}^{d}}J(z)e^{a^{\prime}\left\langle z\right\rangle}\,\mathrm{d}z<\infty.$
(Let us point out that this is where the remark at the end of Theorem 1.1 comes from). Then
$\begin{split}L^{*}_{\varepsilon}e^{a\left\langle x\right\rangle}=\frac{1}{\varepsilon^{2}}\int_{\mathbb{R}^{d}}J_{\varepsilon}(y)(\varphi(x+y)-\varphi(x))\,\mathrm{d}y-x\nabla\varphi&\leq C_{2}e^{a\left\langle x\right\rangle}-a\frac{|x|^{2}}{\left\langle x\right\rangle}e^{a\left\langle x\right\rangle}\leq C-\lambda e^{a\left\langle x\right\rangle}\end{split}$
for some $C,\lambda>0$. ∎
We thus have proved a Lyapunov condition on the generator for $V=\left\langle\cdot\right\rangle^{k}$ and $V=e^{a\left\langle\cdot\right\rangle}$. For these weight functions (and in general for $\mathcal{C}^{2}$ functions), we see that actually Assumption (14) implies Assumption (SLC), which is the statement of the next Theorem.
Theorem 3.5 (Semigroup Lyapunov condition).
Assume $u_{0}$ and $J$ satisfy the hypothesis of Theorem 3.2 (respectively 3.4). For $\varphi=\left\langle x\right\rangle^{k}$ (respectively $\varphi=e^{a\left\langle x\right\rangle}$), the semigroup associated to equation (NLFP) satisfies a Lyapunov condition (SLC) with $V\equiv\varphi$ and the same $C$ and $\lambda$ of Theorem 3.2 (respectively 3.4).
Proof.
We just prove it for $L^{1}_{k}$, being analogous in the other case.
Let $u_{0}$ be in $\mathcal{D}(L_{\varepsilon})$ by standard semigroup theory $u(t)\in\mathcal{C}^{1}([0,\infty);L^{1}_{k})$, and consequently we compute
$\begin{split}\frac{\mathrm{d}}{\mathrm{d}t}\int_{\mathbb{R}^{d}}u(t,\cdot)\varphi(x)\,\mathrm{d}x&=\int_{\mathbb{R}^{d}}\frac{\mathrm{d}}{\mathrm{d}t}u(t,\cdot)\varphi(x)\,\mathrm{d}x\\ &=\int_{\mathbb{R}^{d}}(Lu)\varphi(x)\stackrel{{\scriptstyle\mathclap{\mbox{*}}}}{{=}}\int_{\mathbb{R}^{d}}u(L^{*}\varphi)\leq\int_{\mathbb{R}^{d}}u(-\lambda\varphi+C)\end{split}$
[…]
We then use Gronwall Lemma with the $\mathcal{C}^{1}$ function $f(t):=\int_{\mathbb{R}^{d}}u(t)\varphi$: That is,
$f^{\prime}(t)\leq-\lambda f(t)+C\implies f(t)\leq e^{-\lambda t}f(0)+\frac{C}{\lambda}(1-e^{-\lambda t}).$
Since the $\mathcal{D}(L_{\varepsilon})$ is dense in $L^{1}_{k}$, the claim is true also for any $u_{0}\in L^{1}_{k}$ by classic density argument. ∎
3.3 An $L^{\infty}$ version of Berry-Esseen’s central limit theorem
Berry-Esseen-type theorems are quantitative versions of the central limit theorem which provide an estimate on the error between the normal distribution and the scaled law. A result of this kind will be crucial for our proof in Section 3.4 of positivity for the nonlocal Fokker-Planck equation. The first version of this theorem was proved independently by Berry (1941) and Esseen (1942) for real, centered i.i.d. random variables $X_{1},\dots X_{n}$ with ${\textrm{Var}}\,(X_{1})=\sigma^{2}$ and $\mathbb{E}(|X_{1}|^{3})=\rho$, obtaining a result of the type
$\sup_{x\in\mathbb{R}}|F_{n}(x)-\Phi(x)|\leq\frac{C\rho}{\sqrt{n}\sigma^{3}}$
where $\Phi$ is the cumulative distribution function (c.d.f.) of the normal distribution, $F_{n}$ is the c.d.f. of $\bar{S}_{n}=\frac{1}{\sigma\sqrt{n}}\sum_{i=1}^{n}X_{i}$, and $C$ is a universal constant depending on the dimension $d$. Several generalizations have been developed, proving the same kind of results for not i.i.d. random vectors in $\mathbb{R}^{d}$, as well as improving the precision of the value of $C$. Recent versions can be found in Bentkus (2005) and Raič (2019).
The majority of this kind of results are “weak” results in the sense that they deal with the cumulative distribution function. For our results we would like to work with the density instead. An important result in this direction is Lions and Toscani (1995) in which the authors provide a stronger version of the CLT on the densities, and later in Goudon et al. (2002), Carlen et al. (2010), Hauray and Mischler (2014) . For this paper we need an estimate on the densities of independent but not identically distributed random vectors in $\mathbb{R}^{d}$, so we will extend the proofs of the previous cited papers to our case. Following the ideas in Goudon et al. (2002), we weaken the hypothesis of finiteness of the third moment. We require only a “bit more” than finite variance, paying a price on the speed of convergence to zero of the error.
Let $f:\mathbb{R}^{d}\to[0,\infty)$ be a probability density function such that there exists $s\in(0,1]$ with
$\int_{\mathbb{R}^{d}}xf(x)\,\mathrm{d}x=0,\quad\int_{\mathbb{R}^{d}}x_{i}x_{j}f(x)\,\mathrm{d}x=\delta_{ij},\quad\int_{\mathbb{R}^{d}}|x|^{2+s}f(x)\,\mathrm{d}x:=\rho_{2+s}<\infty.$ (17)
Notice that this normalisation of moments of order $2$ is different from (1) by a factor of $2$. We prefer the present normalisation in this section, which is a natural one in a probabilistic setting and is the same one used in most statements of Berry-Esseen’s theorem. By a simple scaling, it is straightforward to apply the result if one assumes the normalisation (1) instead.
The main result of this section is the following $L^{\infty}$ version of Berry-Esseen’s Theorem for independent, nonidentically distributed random vectors:
Theorem 3.6.
Let $p\in(1,\infty)$ and $f\in L^{1}(\mathbb{R}^{d})\cap L^{p}(\mathbb{R}^{d})$ satisfying (17). Let $(\sigma_{i})_{i\geq 1}$ be a sequence of positive numbers, and define for $n\geq 1$,
$f^{\sigma_{1},\dots\sigma_{n}}(x):=f_{\sigma_{1}}*f_{\sigma_{2}}*\cdots*f_{\sigma_{n}}(x)$
$\mathfrak{f}_{n}(x):=(\bar{\sigma}^{2}n)^{d/2}f^{\sigma_{1},\dots\sigma_{n}}(\sqrt{n}\bar{\sigma}x),$
with $\bar{\sigma}^{2}=\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}^{2}$. Suppose, in addition, that there exist $l,L>0$ such that
$0<l\leq\sigma_{i}\leq L\qquad i=1,\dots,n$
Then, there exist an integer $N=N(p)$ and a constant $C_{BE}$ depending on $\frac{L}{l},p,d,\rho_{2+s},\left\|f\right\|_{L^{p}}$ such that $\mathfrak{f}_{n}\in L^{\infty}$ for all $n\geq N$ and
$\left\|\mathfrak{f}_{n}-G\right\|_{L^{\infty}}\leq\frac{C_{BE}}{n^{s/2}}\quad\text{ for every $n\geq N$,}$
where $G$ is the density of the standard normalized Gaussian.
In order to prove this we first present some preliminary results. The first one is a generalization of the identity $a^{n}-b^{n}=\sum_{i=1}^{n}(a-b)a^{i-1}b^{n-i}$, that can be easily proven by induction:
Lemma 3.7.
For real $a_{1},\dots,a_{n}$ and $b_{1},\dots b_{n}$, for any $n\geq 1$
$\prod_{i=1}^{n}a_{i}-\prod_{i=1}^{n}b_{i}=\sum_{i=1}^{n}\Big[(a_{i}-b_{i})\big(\prod_{j=1}^{i-1}b_{j}\big)\big(\prod_{j=i+1}^{n}a_{j}\big)\Big]$
The second one is a Hölder regularity result for the Fourier transform.
Lemma 3.8.
Let $f:\mathbb{R}^{d}\to\mathbb{R}$ be a probability distribution such that for some $s\in(0,1)$,
$\int_{\mathbb{R}^{d}}f(x)|x|^{2+s}\,\mathrm{d}x=\rho_{2+s}<+\infty.$
Then $\hat{f}\in\mathcal{C}^{2,s}$; that is, $\hat{f}\in\mathcal{C}^{2}$ and for every $x,y\in\mathbb{R}^{d}$
$\max_{|\alpha|=2}|D^{\alpha}\hat{f}(x)-D^{\alpha}\hat{f}(y)|\leq 4\rho_{2+s}|x-y|^{s}.$
Proof.
The $\mathcal{C}^{2}$-regularity is immediate from Fourier transform properties. We are left to show the Hölder-continuity of the derivatives of order $2$, namely, for every multi-index $\alpha$ such that $|\alpha|=2$,
$\begin{split}&\left|D^{\alpha}\hat{f}(\xi)-D^{\alpha}\hat{f}(\eta)\right|=\left|-\widehat{x^{\alpha}f}(\xi)+\widehat{x^{\alpha}f}(\eta)\right|=\left|\int_{\mathbb{R}^{d}}\Big(e^{-ix\eta}-e^{-ix\xi})x^{\alpha}f(x)\,\mathrm{d}x\right|\\ &\leq\int_{|x|\leq R}|e^{-ix\xi}-e^{-ix\eta}||x|^{2}|f(x)\,\mathrm{d}x+\int_{|x|>R}\left|e^{-ix\xi}-e^{-ix\eta}\right||x|^{2}|f(x)\,\mathrm{d}x\\ &\leq 2|\xi-\eta|\int_{|x|\leq R}|x|^{2+s}|x|^{1-s}f(x)\,\mathrm{d}x+2R^{-s}\int_{|x|>R}|x|^{2+s}f(x)\,\mathrm{d}x\\ &\leq 2\rho_{2+s}(|\xi-\eta|R^{1-s}+R^{-s}).\end{split}$
Choosing $R=|\xi-\eta|^{-1}$ we conclude the proof. ∎
The last ingredient we need is the following technical Lemma (that is basically (Hauray and Mischler, 2014, lemma 4.8) with a weaker assumption on moments. We denote by $B_{\delta}(x_{0})$, the ball in $\mathbb{R}^{d}$ centered at $x_{0}$ with radius $\delta$. When the ball is centered at 0, we use $B_{\delta}$ as a shorthand.
Lemma 3.9.
Let $f$ be a probability distribution on $\mathbb{R}^{d}$ satisfying (17). Then
  1. (i)
    There exist $\delta\in(0,1)$ depending on $\rho_{2+s}$ such that
    $\forall\xi\in B_{\delta}\quad|\hat{f}(\xi)|\leq e^{-\frac{|\xi|^{2}}{4}}$
  2. (ii)
    Assume additionally that $f\in L^{p}$, with $p\in(1,\infty]$. For any $\delta>0$ there exists $\kappa=\kappa(\rho_{2+s},\delta,\left\|f\right\|_{L^{p}})\in(0,1)$ such that sup_|ξ|δ|^f(ξ)|κ(δ).
Proof.
  1. (i)
    For $\hat{f}:\mathbb{R}^{d}\to\mathbb{R}$ and for $\xi$ define $h:[0,|\xi|]\to\mathbb{R}$ such that $h(t):=\hat{f}\Big(t\frac{\xi}{|\xi|}\Big)$. Applying twice the fundamental theorem of calculus to $h$ we write
    $\hat{f}(\xi)=h(|\xi|)=1-\frac{|\xi|^{2}}{2}+\int_{0}^{|\xi|}(h^{\prime\prime}(z)-h^{\prime\prime}(0))(z-|\xi|)\,\mathrm{d}z$
    where
    $\begin{split}h^{\prime\prime}(z)=\Big(\frac{\xi}{|\xi|}\Big)^{T}H_{\hat{f}}\Big(z\frac{\xi}{|\xi|}\Big)\Big(\frac{\xi}{|\xi|}\Big).\end{split}$
    Hence, applying Lemma 3.8, we can bound the reminder
    $\begin{split}&\left|\int_{0}^{|\xi|}\Big(\frac{\xi}{|\xi|}\Big)^{T}\Big(H_{\hat{f}}\Big(z\frac{\xi}{|\xi|}\Big)-H_{\hat{f}}(0)\Big)\Big(\frac{\xi}{|\xi|}\Big)(z-|\xi|)\,\mathrm{d}z\right|\\ &\leq\int_{0}^{|\xi|}\left|\!\left|\!\left|(H_{\hat{f}}\Big(z\frac{\xi}{|\xi|}\Big)-H_{\hat{f}}(0)\right|\!\right|\!\right||z-|\xi||\,\mathrm{d}z\\ &\leq 4\rho_{2+s}\int_{0}^{|\xi|}|z|^{s}|z-|\xi||\,\mathrm{d}z=\frac{4\rho_{2+s}|\xi|^{2+s}}{(s+1)(s+2)},\end{split}$
    where $H_{f}(x)$ is the Hessian matrix of $f$ applied at point $x$ and $\left|\!\left|\!\left|\cdot\right|\!\right|\!\right|$ is the spectral norm which is always less or equal than the $\max$- norm used in Lemma 3.8.
    Thus, for small $|\xi|<\delta$ which explicitly depends only on $s$ and $\rho_{2+s}$, we have |^f(ξ)|≤|1-38|ξ|^2|≤e^-|ξ|24.
  2. (ii)
    This statement is a direct consequence of (Carlen et al., 2010, Prop. 26 (iii)), where the assumption of finiteness of the entropy is ensured by our assumption that $f\in L^{1}\cap L^{p}$ and has some finite moment. ∎
We can now present the proof of our version of Berry-Esseen theorem, that follows the ideas presented in (Goudon et al., 2002, Thm. 1) and (Hauray and Mischler, 2014, Thm. 4.6).
Proof of Theorem 3.6.
(NOT THE UPDATED VERSION: $l/L$) First, we notice that $\mathfrak{f}_{n}$ satisfies the hypotheses of Lemma 3.9 for every $n$, and, in particular
$\int_{\mathbb{R}^{d}}\mathfrak{f}_{n}(x)x_{i}x_{k}\,\mathrm{d}x=\delta_{ij}.$
We can rewrite
$\begin{split}{\hat{\mathfrak{f}}_{n}}(\xi)&=\int_{\mathbb{R}^{d}}e^{-ix\xi}(\bar{\sigma}^{2}n)^{d/2}f^{\sigma_{1}\dots,\sigma_{n}}(\sqrt{n}\bar{\sigma}x)\,\mathrm{d}x=\int_{\mathbb{R}^{d}}e^{-iy\frac{\xi}{\sqrt{n}\bar{\sigma}}}f_{\sigma_{1}}*\cdots*f_{\sigma_{n}}(y)\,\mathrm{d}x\\ &=\mathcal{F}(f_{\sigma_{1}}*\cdots*f_{\sigma_{n}})\Big(\frac{\xi}{\sqrt{n}\bar{\sigma}}\Big)=\prod_{j=1}^{n}\hat{f}_{\sigma_{j}}\Big(\frac{\xi}{\sqrt{n}\bar{\sigma}}\Big)=\prod_{j=1}^{n}\int_{\mathbb{R}^{d}}e^{-ix\frac{\xi}{\sqrt{n}\bar{\sigma}}}\sigma_{j}^{-d}f\Big(\frac{x}{\sigma_{j}}\Big)\,\mathrm{d}x\\ &=\prod_{j=1}^{n}\int_{\mathbb{R}^{d}}e^{-iy\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}}f(y)\,\mathrm{d}y=\prod_{j=1}^{n}\hat{f}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big),\end{split}$
and, since $G$ is the standard gaussian, its Fourier transform satisfies
$\hat{G}(\xi)=\prod_{j=1}^{n}\hat{G}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big).$
Without loss of generality, we can assume $p\leq 2$. Hausdorff-Young’s inequality then implies that $\hat{f}$ is in $L^{p^{\prime}}\cap L^{\infty}$ with $p^{\prime}\in[2,\infty)$. Using Hölder’s inequality,
$\left\|\prod_{j=1}^{n}\hat{f}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big)\right\|_{L^{1}}\leq\prod_{j=1}^{n}\left\|\hat{f}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big)\right\|_{L^{n}}=\prod_{j=1}^{n}\Big(\frac{\bar{\sigma}\sqrt{n}}{\sigma_{j}}\Big)^{d/n}\left\|\hat{f}\right\|_{L^{n}}\leq n^{d/2}\Big(\frac{L}{l}\Big)^{d}\left\|\hat{f}\right\|_{L^{n}}^{n},$
which is finite for $n\geq p^{\prime}$, i.e. $\hat{\mathfrak{f}}_{n}$ is in $L^{1}$ for any $n\geq p^{\prime}$. We can thus control the $L^{\infty}$ norm with $L^{1}$ norm of the Fourier transform, since
$\left|\mathfrak{f}_{n}(x)-G(x)\right|=(2\pi)^{-d}\left|\int_{\mathbb{R}^{d}}(\hat{\mathfrak{f}}_{n}(\xi)-\hat{G}(\xi))e^{i\xi x}\,\mathrm{d}\xi\right|\leq(2\pi)^{-d}\int_{\mathbb{R}^{d}}\left|\hat{\mathfrak{f}}_{n}-\hat{G}\right|\,\mathrm{d}\xi$
Let $\delta$ be given by Lemma 3.9 $(i)$ for $\hat{f}$ and define $B_{\delta}=B_{\delta}(n,l,L):=\{\xi:\left|\xi\right|<\sqrt{n}\frac{L}{l}\delta\}$. Following known techniques, we split the integral into high and low frequencies
$\left\|\mathfrak{f}_{n}-G\right\|_{L^{\infty}}\leq\int_{\xi\in B^{c}_{\delta,n}}|\hat{\mathfrak{f}}_{n}|\,\mathrm{d}\xi+\int_{\xi\in B^{c}_{\delta,n}}|\hat{G}|\,\mathrm{d}\xi+\int_{\xi\in B_{\delta,n}}|\hat{\mathfrak{f}}_{n}-\hat{G}|\,\mathrm{d}\xi:=T_{1}+T_{2}+T_{3}$
The first two terms can be bounded in a similar way. For $T_{1}$, and without loss of generality, assume $p^{\prime}$ is an integer (otherwise, we can work with $\left\lceil p^{\prime}\right\rceil$, the smallest integer larger or equal than $p^{\prime}$, since $\hat{f}$ is also in $L^{\left\lceil p^{\prime}\right\rceil}$). Then
$\int_{B_{\delta}^{c}}\prod_{j=1}^{n}\left|\hat{f}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big)\right|\,\mathrm{d}\xi=n^{d/2}\int_{|\eta|\geq\frac{L}{l}\delta}\prod_{j=1}^{n}\left|\hat{f}\Big(\frac{\sigma_{j}}{\bar{\sigma}}\eta\Big)\right|\,\mathrm{d}\eta\\ \leq n^{d/2}\sup_{|\eta|\geq\frac{L}{l}\delta}\left|\hat{f}\Big(\frac{\sigma_{p^{\prime}+1}}{\bar{\sigma}}\eta\Big)\right|\cdots\sup_{|\eta|\geq\frac{L}{l}\delta}\left|\hat{f}\Big(\frac{\sigma_{n}}{\bar{\sigma}}\eta\Big)\right|\int_{|\eta|\geq\frac{L}{l}\delta}\left|\hat{f}\Big(\frac{\sigma_{1}}{\bar{\sigma}}\eta\Big)\right|\cdots\left|\hat{f}\Big(\frac{\sigma_{p^{\prime}}}{\bar{\sigma}}\eta\Big)\right|\,\mathrm{d}\eta.$ (18)
For each $j=p^{\prime}+1,\dots,n$
$\sup_{|\eta|\geq\frac{L}{l}\delta}\left|\hat{f}\Big(\frac{\sigma_{j}}{\bar{\sigma}}\eta\Big)\right|\leq\sup_{|\eta|\geq\frac{\bar{\sigma}}{\sigma_{i}}\delta}\left|\hat{f}\Big(\frac{\sigma_{j}}{\bar{\sigma}}\eta\Big)\right|=\sup_{|\eta^{\prime}|\geq\delta}|\hat{f}(\eta^{\prime})|\leq\kappa(\delta)$ (19)
with $\kappa(\delta)\in(0,1)$, coming from Lemma 3.9 $(ii)$. The factors inside the integral in Equation (18) can be bounded by using Hölder’s inequality: we rewrite
$\begin{split}\int_{|\eta|\geq\frac{L}{l}\delta}&\left|\hat{f}\Big(\frac{\sigma_{1}}{\bar{\sigma}}\eta\Big)\right|\cdots\left|\hat{f}\Big(\frac{\sigma_{p^{\prime}}}{\bar{\sigma}}\eta\Big)\right|\,\mathrm{d}\eta\\ &\leq\bigg(\int_{\mathbb{R}^{d}}\left|\hat{f}\Big(\frac{\sigma_{1}}{\bar{\sigma}}\eta\Big)\,\mathrm{d}\eta\right|^{p^{\prime}}\bigg)^{1/p^{\prime}}\cdots\bigg(\int_{\mathbb{R}^{d}}\left|\hat{f}\Big(\frac{\sigma_{p^{\prime}}}{\bar{\sigma}}\eta\Big)\,\mathrm{d}\eta\right|^{p^{\prime}}\bigg)^{1/p^{\prime}}\\ &=\Big(\frac{\sigma_{1}}{\bar{\sigma}}\Big)^{d/p^{\prime}}\|\hat{f}\|_{p^{\prime}}\cdots\Big(\frac{\sigma_{p^{\prime}}}{\bar{\sigma}}\Big)^{d/p^{\prime}}\|\hat{f}\|_{p^{\prime}}\leq\Big(\frac{L}{l}\Big)^{d}\|\hat{f}\|_{p^{\prime}}^{p^{\prime}}\leq\Big(\frac{L}{l}\Big)^{d}C_{p}\|f\|_{p}^{p^{\prime}},\end{split}$ (20)
after first applying Hölder’s inequality and then Hausdorff-Young’s inequality ($C_{p}$ is the Hausdorff-Young constant). Hence, plugging (19) and (20) into equation (18) we obtain
$T_{1}\leq C_{p}\Big(\sqrt{n}\frac{L}{l}\Big)^{d}\kappa(\delta)^{n-p^{\prime}}\|f\|_{p}^{p^{\prime}}.$ (21)
We proceed analogously for $T_{2}$, obtaining a very similar bound also for $T_{2}$. Since $\kappa(\delta)<1$, $\kappa(\delta)^{n-p^{\prime}}$ decays very fast in $n$, thus for $n$ large enough (depending on $p$), there exists $C_{1}>0$ such that
$T_{1}+T_{2}\leq\frac{C_{1}}{n^{s/2}}$ (22)
Now we focus our attention on $T_{3}$. Using the factorization in Lemma 3.7 we obtain
$\left|\hat{\mathfrak{f}}_{n}(\xi)-\hat{G}(\xi)\right|=\prod_{j=1}^{n}\hat{f}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big)-\prod_{i=1}^{n}\hat{G}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big)\\ =\sum_{j=1}^{n}\left|\hat{f}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big)-\hat{G}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big)\right|\left|\prod_{k=1}^{j-1}\hat{f}\Big(\frac{\xi\sigma_{k}}{\sqrt{n}\bar{\sigma}}\Big)\right|\left|\prod_{k=j+1}^{n}\hat{G}\Big(\frac{\xi\sigma_{k}}{\sqrt{n}\bar{\sigma}}\Big)\right|.$
From Lemma 3.9 $(i)$ it follows that for $\xi\in B_{\delta}$
$\left|\hat{f}\Big(\frac{\xi\sigma_{k}}{\sqrt{n}\bar{\sigma}}\Big)\right|\leq e^{-\frac{|\xi|^{2}}{4n}\frac{\sigma_{k}^{2}}{\bar{\sigma}^{2}}}\leq e^{-\frac{|\xi|^{2}}{4n}(\frac{l}{L})^{2}}.$
and
$\left|\hat{G}\Big({\frac{\xi\sigma_{k}}{\sqrt{n}\bar{\sigma}}}\Big)\right|=(2\pi)^{-d/2}e^{-\frac{|\xi|^{2}}{2n}\frac{\sigma_{k}^{2}}{\bar{\sigma}^{2}}}\leq e^{-\frac{|\xi|^{2}}{2n}(\frac{l}{L})^{2}}.$
Then
$\begin{split}\frac{\left|\hat{\mathfrak{f}}_{n}(\xi)-\hat{G}(\xi)\right|}{|\xi|^{2+s}}&=\frac{1}{\left|\xi\right|^{2+s}}\Big[\sum_{j=1}^{n}\left|\hat{f}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big)-\hat{G}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big)\right|\left|\prod_{k=1}^{j-1}\hat{f}\Big(\frac{\xi\sigma_{k}}{\sqrt{n}\bar{\sigma}}\Big)\right|\left|\prod_{k=j+1}^{n}\hat{G}\Big(\frac{\xi\sigma_{k}}{\sqrt{n}\bar{\sigma}}\Big)\right|\Big]\\ &\leq\frac{1}{\left|\xi\right|^{2+s}}\sum_{j=1}^{n}\left|\hat{f}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big)-\hat{G}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big)\right|e^{-\frac{|\xi|^{2}}{4n}(j-1)(\frac{l}{L})^{2}}e^{-\frac{|\xi|^{2}}{2n}(n-j)(\frac{l}{L})^{2}}\\ &\leq\frac{1}{\left|\xi\right|^{2+s}}\sum_{j=1}^{n}\left|\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\right|^{2+s}\frac{\left|\hat{f}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big)-\hat{G}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big)\right|}{\left|\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\right|^{2+s}}e^{-\frac{|\xi|^{2}}{4}\frac{n-1}{n}(\frac{l}{L})^{2}}\\ &\leq\frac{1}{n^{\frac{2+s}{2}}}\Big(\frac{L}{l}\Big)^{2+s}n\sup_{\eta^{\prime}}\frac{|\hat{f}(\eta^{\prime})-\hat{G}(\eta^{\prime})|}{\left|\eta^{\prime}\right|^{2+s}}e^{-\frac{\left|\xi\right|^{2}}{8}(\frac{l}{L})^{2}}.\end{split}$
Since $f$ and $G$ are both centered and have the same covariance matrix, in the same spirit of the proof of Lemma 3.9 $(i)$, we obtain
$\sup_{\eta^{\prime}}\frac{|\hat{f}(\eta^{\prime})-\hat{G}(\eta^{\prime})|}{\left|\eta^{\prime}\right|^{2+s}}\leq\sup_{\eta^{\prime}}\left(\frac{|\hat{f}(\eta^{\prime})-\hat{f}(0)|}{\left|\eta^{\prime}\right|^{2+s}}+\frac{|\hat{G}(\eta^{\prime})-\hat{G}(0)|}{\left|\eta^{\prime}\right|^{2+s}}\right)\leq C_{1}(\rho_{2+s}+M_{2+s}(G)).$
where $M_{2+s}(G)=\int|x|^{2+s}G(x)\,\mathrm{d}x$. This leads to the final estimate on the third term (with the “correct” rate of decay), that is
$\begin{split}T_{3}&\leq\frac{C_{1}}{n^{s/2}}\Big(\frac{L}{l}\Big)^{2+s}(\rho_{2+s}+M_{2+s}(G))\int_{\mathbb{R}^{d}}e^{-\frac{\left|\xi\right|^{2}}{8}(\frac{l}{L})^{2}}|\xi|^{2+s}\,\mathrm{d}\xi\\ &:=\frac{1}{n^{s/2}}\Big(\frac{L}{l}\Big)^{2+s}(\rho_{2+s}+M_{2+s}(G))C_{\frac{L}{l},d}.\end{split}$
Taking into account (22) and adding $T_{3}$, the proof is completed. ∎
3.4 Positivity of the solution
In order to show Harris’s positivity condition (Hypothesis (HC)) we are in fact able to prove a stronger result, namely a local positivity condition at any positive time $t>0$:
Theorem 3.10 (Lower bound).
Let $J\colon\mathbb{R}^{d}\to[0,\infty)$ satisfy (1), (3), and $J\in L^{p}$ for some $p\in(1,\infty].$ Then, the solution $u$ of equation (NLFP) with a probability initial condition $u_{0}$ is uniformly positive on compact sets for all $0<\varepsilon\leq 1$. That is: for every $0<\varepsilon\leq 1$, $t>0$, $R_{1}R_{2}>0$, there exists an $\varepsilon$-independent constant $\alpha,$ such that the solution of (NLFP) with initial data $u_{0}\in L^{1}_{k}$ is such that
$u(t,x)\geq\alpha\int_{B_{R_{2}}}u_{0}\qquad\text{for }x\in B_{R_{1}}.$
This is a nontrivial estimate since we seek the independence of $\alpha$ from $\varepsilon$; in particular, we seek positivity that holds uniformly in the limit $\varepsilon\rightarrow 0$.
Our proof of this estimate is based on the Wild sum representation of the solution from Section 2.3 and on the version of Berry-Esseen’s theorem we proved in Section 3.3. We distinguish two cases which are described by the two following Lemmas:
Lemma 3.11.
Let $J$ satisfy (1) and (3). Then, for any $t,\eta>0$, there exists $\varepsilon_{0}>0$ and a ($\varepsilon$-independent) constant $A>0$ such that
$J_{\varepsilon}^{t_{1},\dots,t_{n}}(x)\geq A\qquad\text{ for all }x\in B_{\eta}$
for all $\varepsilon<\varepsilon_{0}$ and for any $t_{1},\dots,t_{n}$ with $t\geq t_{1}\geq\dots t_{n}\geq 0$ and $n$ such that
$\frac{t}{\varepsilon^{2}}\leq n\leq 2\frac{t}{\varepsilon^{2}}.$
Proof.
Define the independent random vectors $X_{i}\sim J_{\varepsilon e^{t_{i}}}$ for which
$\int_{\mathbb{R}^{d}}xJ_{\varepsilon e^{t_{i}}}\,\mathrm{d}x=0\qquad\int_{\mathbb{R}^{d}}x_{j}x_{k}J_{\varepsilon e^{t_{i}}}(x)\,\mathrm{d}x=2\varepsilon^{2}e^{2t_{i}}\delta_{jk}.$
This implies that $S_{n}=\sum_{i=1}^{n}X_{i}\sim J_{\varepsilon}^{t_{1},\dots,t_{n}}$ and
$\int_{\mathbb{R}^{d}}xJ_{\varepsilon}^{t_{1},\dots,t_{n}}(x)\,\mathrm{d}x=0\qquad\int_{\mathbb{R}^{d}}x_{j}x_{k}J_{\varepsilon}^{t_{1},\dots,t_{n}}(x)\,\mathrm{d}x=2\varepsilon^{2}n\bar{\tau}^{2}\delta_{jk},$
where $\bar{\tau}^{2}:=\frac{1}{n}\sum_{i=1}^{n}e^{2t_{i}}$ and $1\leq\bar{\tau}^{2}\leq e^{2t}$. Notice that
$l:=\varepsilon\leq\varepsilon e^{t_{n}}\leq\dots\leq\varepsilon e^{t_{1}}\leq\varepsilon e^{t}:=L$
and thus $\frac{L}{l}=e^{t}.$ Let $N>0$ be the threshold for which the Berry-Esseen estimate 3.6 holds, and define $\varepsilon_{1}=\sqrt{\frac{t}{N}}$. Let $\varepsilon_{2}>0$ be such that
$\Big[G\Big(\frac{\eta}{\sqrt{2t}}\Big)-\frac{C_{BE}}{t^{s/2}}\varepsilon_{2}^{s}\Big]\geq\frac{1}{2}G\Big(\frac{\eta}{\sqrt{2t}}\Big).$
Take $\varepsilon_{0}:=\min\{\varepsilon_{1},\varepsilon_{2}\}$. For all $\varepsilon<\varepsilon_{0}$, we can apply Theorem 3.6, since, by hypothesis, $n\geq\frac{t}{\varepsilon^{2}}\geq\frac{t}{\varepsilon_{1}^{2}}=N$. Hence, since $G$ is radially decreasing,
$\begin{split}J_{\varepsilon}^{t_{1},\dots,t_{n}}(x)&\geq(2\varepsilon^{2}n\bar{\tau}^{2})^{-d/2}\Big[G\Big(\frac{x}{\sqrt{2n}\varepsilon\bar{\tau}}\Big)-\frac{C_{BE}}{n^{s/2}}\Big]\\ &\geq(2\varepsilon^{2}n\bar{\tau}^{2})^{-d/2}\Big[G\Big(\frac{\eta}{\sqrt{2t}\bar{\tau}}\Big)-\frac{C_{BE}}{t^{s/2}}\varepsilon^{s}\Big]\\ &\geq(2\varepsilon^{2}n\bar{\tau}^{2})^{-d/2}\Big[G\Big(\frac{\eta}{\sqrt{2t}\bar{\tau}}\Big)-\frac{C_{BE}}{t^{s/2}}\varepsilon_{2}^{s}\Big]\\ &\geq(4te^{2t})^{-d/2}\frac{1}{2}G\Big(\frac{\eta}{\sqrt{2t}}\Big):=A.\qed\end{split}$ (23)
If $\varepsilon$ has a positive lower bound $\varepsilon_{0}$, then we prove the following Lemma in the same spirit:
Lemma 3.12.
Let $J$ satisfies (1) and (3). Take $\varepsilon_{0}\in(0,1]$, $t>0$ and $\eta>0$. Then, there exist an explicit $N\in\mathbb{N},B>0$ such that for all $n\geq N$ and $\varepsilon\in[\varepsilon_{0},1]$,
$J_{\varepsilon}^{t_{1},\dots,t_{n}}(x)\geq\frac{B}{n^{\frac{d}{2}}}=:B_{n}\qquad\text{for all }x\in B_{\eta}$
where $B$ depends on $\varepsilon_{0}$, $t$, $d$, $\eta$ and the Berry-Esseen constant.
Proof.
Let $N=\max\{N_{1},N_{2}\}$, where $N_{1}$ is the threshold for which estimate in Theorem 3.6 holds, and $N_{2}$ is such that
$G\Big(\frac{\eta}{\sqrt{2}\varepsilon_{0}}\Big)-\frac{C_{BE}}{N_{2}^{s/2}}\geq\frac{1}{2}G\Big(\frac{\eta}{\sqrt{2}\varepsilon_{0}}\Big).$
Then for $n\geq N$, applying Berry - Esseen CLT, we obtain
$\begin{split}J_{\varepsilon}^{t_{1},\dots,t_{n}}(x)&\geq(2\varepsilon^{2}n\bar{\tau}^{2})^{-d/2}\Big[G\Big(\frac{x}{\sqrt{2n}\varepsilon\bar{\tau}}\Big)-\frac{C_{BE}}{n^{s/2}}\Big]\\ &\geq\frac{1}{n^{\frac{d}{2}}}\Big(\frac{1}{\sqrt{2}e^{t}}\Big)^{d}\Big[G\Big(\frac{\eta}{\sqrt{2}\varepsilon_{0}}\Big)-\frac{C_{BE}}{N_{2}^{s/2}}\Big]\\ &\geq\frac{1}{n^{\frac{d}{2}}}\Big(\frac{1}{\sqrt{2}e^{t}}\Big)^{d}\frac{1}{2}G\Big(\frac{\eta}{\sqrt{2}\varepsilon_{0}}\Big)=:\frac{B}{n^{\frac{d}{2}}}.\end{split}$
The last ingredient we need to prove Theorem 3.10 is the following lemma on the partial sums of the densities of Poisson-distributed random variables.
Lemma 3.13.
There exists an explicit constant $C_{L}$ such that
$s_{m}:=e^{-m}\sum_{n=m}^{2m}\frac{m^{n}}{n!}\geq C_{L}\qquad\text{for all $m\geq 1$,}$
and $\lim_{m\to\infty}s_{m}=\frac{1}{2}.$
Proof.
the la (not the updated version) Let $X_{m}$ be a Poisson random variable of parameter $m$. For a fixed $m$, using a known result on bounds for the tail probabilities of a Poisson random variable, we have that
$s_{m}=1-\P(X_{m}<m)-\P(X_{m}\leq 2m+1)\geq\frac{1}{2}-e^{m+1}\Big(\frac{m}{2m+1}\Big)^{2m+1}:=b_{m}\geq b_{1}=:C_{L},$
since $b_{m}$ is increasing for $m\geq 1$. To prove that $s_{m}\to 1/2$, let us now take a random variable $X_{m}$, Poisson-distributed with parameter $m$. Let $Y_{i},i=1,\dots,m$ be Poisson-distributed random variables with parameter $1$ so that $X_{m}=\sum_{i=1}^{m}Y_{i}$. We have
$\begin{split}&e^{-m}\sum_{n=m}^{m}\frac{m^{n}}{n!}=e^{-m}\sum_{n=0}^{2m}\frac{m^{n}}{n!}-e^{-m}\sum_{n=0}^{m-1}\frac{m^{n}}{n!}\\ &=\mathbb{P}\Big(\sum_{i=1}^{m}Y_{i}-m\leq{m}\Big)-\mathbb{P}\Big(\frac{1}{\sqrt{m}}\big({\sum_{i=1}^{m}Y_{i}-m}\big)\leq 0\Big)\ \rightarrow 1-\frac{1}{2}=\frac{1}{2},\end{split}$
thanks to the law of large numbers and the central limit theorem. ∎
Remark 3.14.
For our proofs we will use the above lemma as stated. However, with a very similar proof one can show that for any $r>1$, the lemma still holds with a partial sum from $m$ to $\left\lceil rm\right\rceil$ (instead of $m$ to $2m$), with the same limit of $1/2$.
Proving Theorem 3.10 is now straightforward:
Proof of Theorem 3.10.
Expressing the solution via Wild sums we have
$u(t,x)\geq e^{(d-\frac{1}{\varepsilon^{2}})t}\sum_{n=1}^{\infty}\varepsilon^{-2n}\int_{0}^{t}\dots\int_{0}^{t_{n-1}}J_{\varepsilon}^{t_{1},\dots,t_{n}}*u_{0}(e^{t}x)\,\mathrm{d}t_{n}\dots\,\mathrm{d}t_{1}.$
Set $\eta=e^{t}R_{1}+R_{2}$ and take $\varepsilon_{0}$ from Lemma 3.11. If $\varepsilon\leq\tilde{\varepsilon}_{0}:=\min\{\varepsilon_{0},\sqrt{t}\}$, applying Lemma 3.11 we have
$\begin{split}u(t,x)&\geq e^{(d-\frac{1}{\varepsilon^{2}})t}\sum_{n=1}^{\infty}\varepsilon^{-2n}\int_{0}^{t}\dots\int_{0}^{t_{n-1}}\int_{\mathbb{R}^{d}}J_{\varepsilon}^{t_{1},\dots,t_{n}}(e^{t}x-y)u_{0}(y)\,\mathrm{d}y\,\mathrm{d}t_{n}\dots\,\mathrm{d}t_{1}\\ &\geq e^{(d-\frac{1}{\varepsilon^{2}})t}\sum_{n=\left\lfloor\frac{t}{\varepsilon^{2}}\right\rfloor}^{2\left\lfloor\frac{t}{\varepsilon^{2}}\right\rfloor}\varepsilon^{-2n}\int_{0}^{t}\dots\int_{0}^{t_{n-1}}\int_{B_{R_{2}}}J_{\varepsilon}^{t_{1},\dots,t_{n}}(e^{t}x-y)u_{0}(y)\,\mathrm{d}y\,\mathrm{d}t_{n}\dots\,\mathrm{d}t_{1}\\ &\geq Ae^{dt}e^{-\frac{t}{\varepsilon^{2}}}\sum_{n=\left\lfloor\frac{t}{\varepsilon^{2}}\right\rfloor}^{2\left\lfloor\frac{t}{\varepsilon^{2}}\right\rfloor}\Big(\frac{t}{\varepsilon^{2}}\Big)^{n}\frac{1}{n!}\int_{B_{R_{2}}}u_{0}(y)\,\mathrm{d}y\\ &\geq\Big(Ae^{\,\mathrm{d}t}e^{-(\frac{t}{\varepsilon^{2}}-\left\lfloor\frac{t}{\varepsilon^{2}}\right\rfloor)}\int_{B_{R_{2}}}u_{0}(y)\,\mathrm{d}y\Big)e^{-\left\lfloor\frac{t}{\varepsilon^{2}}\right\rfloor}\sum_{n=\left\lfloor\frac{t}{\varepsilon^{2}}\right\rfloor}^{2\left\lfloor\frac{t}{\varepsilon^{2}}\right\rfloor}\left\lfloor\frac{t}{\varepsilon^{2}}\right\rfloor^{n}\frac{1}{n!}\\ &\geq\Big(Ae^{dt-1}\int_{B_{R_{2}}}u_{0}(y)\,\mathrm{d}y\Big)e^{-m}\sum_{n=m}^{2m}\frac{m^{n}}{n!}\geq AC_{L}e^{dt-1}\int_{B_{R_{2}}}u_{0}(y)\,\mathrm{d}y\end{split}$
with $m=\left\lfloor\frac{t}{\varepsilon^{2}}\right\rfloor\geq 1$ and $C_{L}$ the constant found in Lemma 3.13. Notice that we have used $\tilde{\varepsilon}_{0}\leq\varepsilon_{0}$ in order to apply Lemma 3.11, and also $\tilde{\varepsilon}_{0}\leq\sqrt{t}$ to make sure that the bound we obtain is not trivial (that is, to avoid $m=0$).
If $\varepsilon\geq\tilde{\varepsilon}_{0}$ then, applying Lemma 3.12 to $\tilde{\varepsilon}_{0}$, for all $x\in B_{\eta}$ we obtain
$\begin{split}u(t,x)&\geq e^{(d-\frac{1}{\varepsilon^{2}})t}\sum_{n=1}^{\infty}\varepsilon^{-2n}\int_{0}^{t}\dots\int_{0}^{t_{n-1}}\int_{\mathbb{R}^{d}}J_{\varepsilon}^{t_{1},\dots,t_{n}}(e^{t}x-y)u_{0}(y)\,\mathrm{d}y\,\mathrm{d}t_{n}\dots\,\mathrm{d}t_{1}\\ &\geq e^{(d-\frac{1}{\varepsilon^{2}})t}\sum_{n=1}^{\infty}\varepsilon^{-2n}\int_{0}^{t}\dots\int_{0}^{t_{n-1}}\int_{B_{R_{2}}}J_{\varepsilon}^{t_{1},\dots,t_{n}}(e^{t}x-y)u_{0}(y)\,\mathrm{d}y\,\mathrm{d}t_{n}\dots\,\mathrm{d}t_{1}\\ &\geq e^{(d-\frac{1}{\varepsilon^{2}})t}\sum_{n=N}^{\infty}\Big(\frac{t}{\varepsilon^{2}}\Big)^{n}\frac{1}{n!}B_{n}\int_{B_{R_{2}}}u_{0}(y)\,\mathrm{d}y\\ &\geq\frac{B}{N!}e^{(d-\frac{1}{\tilde{\varepsilon}_{0}^{2}})t}\Big(\frac{t}{\varepsilon_{0}^{2}}\Big)^{N}\frac{1}{N^{\frac{d}{2}}}\int_{B_{R_{2}}}u_{0}(y)\,\mathrm{d}y.\end{split}$
We finally take
$\alpha:=\min\Big\{AC_{L}e^{dt-1},\frac{B}{N!}e^{(d-\frac{1}{\varepsilon_{0}^{2}})t}\Big(\frac{t}{\tilde{\varepsilon}_{0}^{2}}\Big)^{N}\frac{1}{N^{\frac{d}{2}}}\Big\}.$
Remark 3.15.
The previous proof of positivity works almost without changes for the equation
$\partial_{t}u=\frac{1}{\varepsilon^{2}}(J_{\varepsilon}*u-u)+\operatorname{div}(\nabla Vu)$
for any, even time-dependent, potential $V:\mathbb{R}^{d}\times\mathbb{R}\to\mathbb{R}$ such that
$\frac{d}{\,\mathrm{d}t}X(t)=\nabla V(X(t),t)$
is a linear ODE. It seems harder to extend straightforwardly the proof to potentials with different expressions.
Proof of the main result
We have now all the ingredients to prove Theorem 1.1.
Proof.
The proof comes straightforwardly thanks to Harris’s Theorem 3.1. From Theorem 3.10, at every time $t>0$ and for every $R_{1},R_{2}>0$, the operator $S_{t}$ satisfies Hypothesis (HC):
$S_{t}u_{0}\geq\alpha\mathds{1}_{B_{R_{1}}}\int_{\mathcal{S}}u_{0}(y)\,\mathrm{d}y$
on the “small set” $\mathcal{S}=B_{R_{2}}$.
(put it twice: once after Lyapunov and once here) Theorem 3.2 gives the confining condition on the generator $L_{\varepsilon}$ in $L^{1}_{\left\langle k\right\rangle}$ and consequently in $L^{1}_{k}$. Let $u_{0}$ be in $\mathcal{D}(L_{\varepsilon})$. By standard semigroup theory $u(t)\in\mathcal{C}^{1}([0,\infty);L^{1}_{k})$, and consequently we compute with $\varphi=\left\langle x\right\rangle^{k}$,
$\begin{split}\frac{\mathrm{d}}{\mathrm{d}t}\int_{\mathbb{R}^{d}}u(t,\cdot)\varphi(x)\,\mathrm{d}x&=\int_{\mathbb{R}^{d}}\frac{\mathrm{d}}{\mathrm{d}t}u(t,\cdot)\varphi(x)\,\mathrm{d}x\\ =&\int_{\mathbb{R}^{d}}(L_{\varepsilon}u)\varphi(x)\stackrel{{\scriptstyle\mathclap{\mbox{*}}}}{{=}}\int_{\mathbb{R}^{d}}u(L_{\varepsilon}^{*}\varphi)\leq\int_{\mathbb{R}^{d}}u(-\lambda\varphi+C)\end{split}$
[…]
$f^{\prime}(t)\leq-\lambda f(t)+C\implies f(t)\leq e^{-\lambda t}f(0)+\frac{C}{\lambda}(1-e^{-\lambda t})$
Since $\mathcal{D}(L_{\varepsilon})$ is dense in $L^{1}_{k}$, the claim is true also for any $u_{0}\in L^{1}_{k}$ by standard density arguments, hence proving the Lyapunov condition on semigroups. Similarly, part $(ii)$ of Theorem 1.1 is an immediate consequence of Theorem 3.10 and Theorem 3.4, proving a semigroup Lyapunov condition in a similar manner with $\varphi=e^{a\left\langle x\right\rangle}.$
Remark 3.16 (On the dependence of the constants).
As mentioned in the introduction, Harris’s theorem provides a constructive method that yields explicit—though generally nonoptimal—constants. Unsurprisingly, the kernel $J$ plays a central role in determining these constants. Indeed, the latter get better the closer $J$ resembles a Gaussian distribution. This is reflected in our use of the Berry–Esseen theorem, which introduces an explicit dependence on both the Berry–Esseen constant and the number $N$ of convolution iterations. We also note that if $J$ were itself a Gaussian, positivity would follow directly and there would be no need to invoke any central limit theorem.
4 Convergence: from nonlocal to local
As a consequence of Theorem 1.1, proved in Section 3, will now prove results on convergence of the nonlocal Fokker-Planck equation (NLFP) to the local one (FP). We present proofs which give a speed of convergence of order $\varepsilon^{2}$, assuming the “symmetry” condition (4) on the kernel $J$. As we mentioned in Remark 1.4, one may make minor modifications in the proofs below in order to avoid condition (4), at the price of obtaining a slower speed of convergence of order $\varepsilon$. Since the proofs are very similar, for the sake of brevity we present only the $\varepsilon^{2}$ case. We will thus assume (4) for the whole section.
We follow a strategy of proof which is common in numerical analysis. We begin by proving suitable consistency and stability results in order to prove convergence for finite times, assuming nice regularity and decay hypotheses.
Proposition 4.1 (Consistency).
Take $k\geq 0$. Assume $J$ satisfies (1), (4), and suppose additionally $J\in L^{1}_{k+d+4}$. Let $v\in\mathcal{C}^{4}\cap L^{1}_{k}$ be a function such that there exist $c>0$ and $K>k+d$ with $|D^{\alpha}v(x)|\leq c\left\langle x\right\rangle^{-K}$ for all multiindices $\alpha$ with $|\alpha|=4$.
Then, there exists a constant $C=C(K,d,k,\left\|J\right\|_{L^{1}_{4+k+d}},c)>0$ such that
$\left\|(L_{\varepsilon}-L_{0})v\right\|_{L^{1}_{k}}\leq C\varepsilon^{2}$
Proof.
To lighten the notation we denote by $\varphi=\left\langle\cdot\right\rangle^{-K}$. We write (understanding that all operators are evaluated at $x\in\mathbb{R}^{d}$)
$L_{\varepsilon}v-L_{0}v=\frac{1}{\varepsilon^{2}}[J_{\varepsilon}*v-v]-\Delta v=\int_{\mathbb{R}^{d}}J(z)[\frac{1}{\varepsilon^{2}}(v(x-\varepsilon z)-v(x))-\Delta_{x}v]\,\mathrm{d}z,$
and we Taylor-expand $v$ around $x$ up to the fourth order:
$v(x-\varepsilon z)=\sum_{|\alpha|\leq 3}D^{\alpha}v(x)\frac{(-\varepsilon z)^{\alpha}}{\alpha!}+\sum_{|\alpha|=4}D^{\alpha}v(\xi_{x,\varepsilon z})\frac{(-\varepsilon z)^{\alpha}}{\alpha!}$
with $\xi_{x,\varepsilon z}\in[x,x-\varepsilon z]$. From assumptions (1) and (4) when integrating in $z$, for $\alpha$ with odd entry $\alpha_{i}$, all terms in the first sum disappear. This leads to
$\begin{split}&\left|L_{\varepsilon}v-L_{0}v\right|=\left|\int_{\mathbb{R}^{d}}J(z)[\frac{1}{\varepsilon^{2}}(v(x-\varepsilon z)-v(x))-\Delta_{x}v]\,\mathrm{d}z\right|\\ &\leq\varepsilon^{2}\int_{\mathbb{R}^{d}}J(z)\sum_{|\alpha|=4}\frac{1}{\alpha!}\left|D^{\alpha}v\right|(\xi_{x,\varepsilon z})|z|^{4}\,\mathrm{d}z\leq c\varepsilon^{2}\int J(z)\varphi(\xi_{x,\varepsilon z})|z|^{4}\,\mathrm{d}z.\end{split}$
Thus, we are left with
$\begin{split}\left\|(L_{\varepsilon}-L_{0})v\right\|_{L^{1}_{k}}&\leq c\varepsilon^{2}\int_{\mathbb{R}^{d}}(1+|x|^{k})\int_{\mathbb{R}^{d}}J(z)\varphi(\xi_{x,\varepsilon z})|z|^{4}\,\mathrm{d}z\,\mathrm{d}x\\ &=c\varepsilon^{2}\int_{\mathbb{R}^{d}}J(z)|z|^{4}\int_{\mathbb{R}^{d}}\varphi(\xi_{x,\varepsilon z})(1+|x|^{k})\,\mathrm{d}x\,\mathrm{d}z.\end{split}$
We split the inner integral:
$\int_{\mathbb{R}^{d}}\varphi(\xi_{x,\varepsilon z})(1+|x|^{k})\,\mathrm{d}x=\int_{|x|\leq 2\varepsilon|z|}\varphi(\xi_{x,\varepsilon z})(1+|x|^{k})\,\mathrm{d}x+\int_{|x|>2\varepsilon|z|}\varphi(\xi_{x,\varepsilon z})(1+|x|^{k})\,\mathrm{d}x.$
The first term can be bounded by
$\int_{|x|\leq 2\varepsilon|z|}\varphi(\xi_{x,\varepsilon z})(1+|x|^{k})\,\mathrm{d}x\leq\varphi(0)(1+|2\varepsilon|^{k}|z|^{k})\left|B_{2\varepsilon|z|}\right|\leq C_{1}\varepsilon^{d}|z|^{d}+C_{2}\varepsilon^{k+d}|z|^{k+d}$
Since in the second integral $\varepsilon|z|<\frac{|x|}{2}$ and $\varphi$ is radial, we can bound
$\begin{split}&\int_{|x|>2\varepsilon|z|}\varphi(\xi_{x,\varepsilon z})(1+|x|^{k})\,\mathrm{d}x\leq\int_{|x|>2\varepsilon|z|}\varphi(|x|-\varepsilon|z|)(1+|x|^{k})\,\mathrm{d}x\\ &\leq\int_{|x|>2\varepsilon|z|}\varphi(x/2)(1+|x|^{k})\,\mathrm{d}x\leq 2^{k+d}\left\|\varphi\right\|_{L^{1}_{k}}\end{split}$
Then
$\begin{split}&\left\|(L_{\varepsilon}-L_{0})v\right\|_{L^{1}_{k}}\leq c\varepsilon^{2}\int_{\mathbb{R}^{d}}J(z)|z|^{4}\big[2^{k+d}\left\|\varphi\right\|_{L^{1}_{k}}+C_{2}\varepsilon^{d}|z|^{d}+C_{3}\varepsilon^{k+d}|z|^{k+d}\big]\,\mathrm{d}z\\ &\leq c\varepsilon^{2}\Big[C_{1}\int_{\mathbb{R}^{d}}J(z)|z|^{4}\,\mathrm{d}z+C_{2}\varepsilon^{d}\int_{\mathbb{R}^{d}}J(z)|z|^{4+d}\,\mathrm{d}z+C_{3}\varepsilon^{k+d}\int_{\mathbb{R}^{d}}J(z)|z|^{4+k+d}\,\mathrm{d}z\Big]\\ &\leq c\varepsilon^{2}\Big(C_{1}\left\|J\right\|_{L^{1}_{4}}+C_{2}\varepsilon^{d}\left\|J\right\|_{L^{1}_{4+d}}+C_{3}\varepsilon^{k+d}\left\|J\right\|_{L^{1}_{4+k+d}}\Big)\leq C\varepsilon^{2}\end{split}$
where $C$ only depends on $c,k,K,d,$, and $\left\|J\right\|_{L^{1}_{4+k+d}}$. ∎
Remark 4.2.
Notice that for any $v$ such that $D^{\alpha}v=0$ with $|\alpha|\leq 4$, like polynomials of order $3$ or lower, the local and the nonlocal operators agree.
We now study the stability of our nonlocal PDE under small perturbations:
Proposition 4.3 (Stability).
Let $u\in\mathcal{C}([0,\infty);L^{1}_{k})$ be the solution of
$\partial_{t}u=L_{\varepsilon}u$
Let $h,v\in\mathcal{C}([0,\infty);L^{1}_{k})$ satisfy in the mild sense
$\partial_{t}v=L_{\varepsilon}v+h(t)$
with the same initial data $u(0)=v(0)=u_{0}.$ Then, there exists constants $\omega>0,M\geq 1$ depending only on $k$ and $d$ such that
$\left\|u(t)-v(t)\right\|_{L^{1}_{k}}\leq M\int_{0}^{t}e^{\omega(t-s)}\left\|h(s)\right\|_{L^{1}_{k}}\,\mathrm{d}s\qquad\text{for all $t\geq 0$.}$
Proof.
$\psi:=v-u$ solves
$\begin{cases}\partial_{t}\psi=L_{\varepsilon}\psi+h(t)\\ \psi(0)=0\end{cases}$
Using Duhamel’s formula we have
$\begin{split}\left\|\psi(t)\right\|_{L^{1}_{k}}&=\left\|\int_{0}^{t}e^{(t-s)L_{\varepsilon}}h(s)\,\mathrm{d}s\right\|_{L_{1}^{k}}\leq M\int_{0}^{t}e^{\omega(t-s)}\left\|h(s)\right\|_{L^{1}_{k}}\,\mathrm{d}s,\end{split}$
since for a semigroup there exist $\omega>0$, $M\geq 1$, such that
$\left\|e^{tL_{\varepsilon}}f\right\|_{L^{1}_{k}}\leq Me^{\omega t}\left\|f\right\|_{L^{1}_{k}}.\qed$
Remark 4.4.
One could improve the growth rate in the above result by a more careful estimate of $e^{(t-s)L_{\varepsilon}}h(s)$. However, since the point of this is to later obtain a local-in-time bound, we give only the above rough estimate.
We leave here a standard result of the decay of the derivatives of the solution of the Fokker-Planck equation.
Lemma 4.5.
Let $u$ be the solution of Fokker-Planck equation with initial data $u_{0}$, $K\geq 0$ and $\alpha$ a multiindex. If $u_{0}$ is such that
$\left|D^{\alpha}u_{0}(x)\right|\leq c\left\langle x\right\rangle^{-K}$
then for every $T^{*}>0$ there exists a constant $C_{T^{*}}$ depending only on $T^{*}$, $K$, $\alpha$ and the dimension $d$ such that
$\left|D^{\alpha}u(x)\right|\leq C_{T^{*}}\left\langle x\right\rangle^{-K}\qquad\text{for all $0\leq t\leq T^{*}$.}$
Sketch of proof.
Since $u(t,x)=e^{dt}v(\frac{1}{2}(e^{2t}-1),xe^{t})$ where $v$ is a solution to the standard heat equation with the same initial data, it is easier to prove the corresponding result for $v$. Since
$v=\Gamma_{t}*u_{0},$
where $\Gamma_{t}$ is the standard heat kernel, passing the derivative to the initial data one easily obtains the result for $v$. Then going back to $u$ one finds a time dependent constant $C_{t}$ such that
$\left|D^{\alpha}u(x)\right|\leq C_{t}\left\langle x\right\rangle^{-K}$
Taking the supremum over $0,T^{*}]$ the proof is complete. ∎
We can therefore prove the following:
Proposition 4.6 (Local-in-time convergence).
Let $J$ satisfy assumptions (1), (4), and assume $J\in L^{1}_{k+d+4}$ $k\geq 0$. Let $u$ be the solution of the nonlocal Fokker-Planck equation (NLFP) and $v$ the solution of the local Fokker-Planck equation (FP) with the same initial data $u_{0}\in L^{1}_{k}\cap\mathcal{C}^{4}$. Suppose that there exists $c>0$ and $K>k+d$ with $|D^{\alpha}u_{0}(x)|\leq c\left\langle x\right\rangle^{-K}$ for all multiindices $\alpha$ with $|\alpha|=4$. Then for all $T^{*}>0$ there exist $C,\omega>0$ such that,
$\left\|u(t)-v(t)\right\|_{L^{1}_{k}}\leq Ce^{\omega T^{*}}\varepsilon^{2}\qquad\text{for $t\in[0,T^{*}]$.}$
Proof.
By Lemma 4.5, the solution $v=v(t,x)$ of (FP) is in $\mathcal{C}^{4}$, for every $t\in[0,T^{*}]$ and $|D^{\alpha}v(t,x)|\leq c\left\langle x\right\rangle^{-K}$ uniformly in $0,T^{*}]$, due to our assumptions on the derivatives of $u_{0}$. Hence, $v$ satisfies the hypotheses of Proposition 4.1, uniformly for all $t\in[0,T^{*}]$. We rewrite
$L_{0}v=L_{\varepsilon}v+(L_{0}v-L_{\varepsilon}v):=L_{\varepsilon}v+h(t).$
Then
$\left\|u(t)-v(t)\right\|_{L^{1}_{k}}\leq Me^{\omega t}\int_{0}^{t}e^{-\omega s}\left\|L_{0}v-L_{\varepsilon}v\right\|_{L^{1}_{k}}\leq M\frac{e^{\omega t}-1}{\omega}C_{1}\varepsilon^{2}\leq Ce^{\omega T^{*}}\varepsilon^{2}$
where we used our results on stability of Proposition 4.3 in the first inequality and the consistency with constant $C_{1}$ of Proposition 4.1 in the second inequality. ∎
Once we have studied the convergence for finite times, we use Theorem 1.1 to gather information on the equilibrium and the behaviour of the solution at large times. As a first consequence of the previous results, we can estimate the speed convergence of the equilibrium $F_{\varepsilon}$ of (NLFP) to the equilibrium $G$ of (FP) as $\varepsilon\to 0$. Indeed, having a spectral gap for an operator (semigroup) provides information of the norm of the resolvent of the latter, assuring that the norm of the resolvent remains bounded by a constant.
Proposition 4.7 (Convergence of the equilibrium).
Let $k>0$ and suppose $J$ satisfies (1) and (4). Additionally assume $J\in L^{1}_{k+d+4}$. Then
$\left\|F_{\varepsilon}-G\right\|_{L^{1}_{k}}\leq C\varepsilon^{2}$
Proof.
Feller-Miyadera-Philips’s general case of Hille-Yosida Theorem for semigroups—see (Engel and Nagel, 2001, Part 2, Thm. 3.8)—states that having an estimate for a strongly continuous semigroup $(S_{t})$ in a Banach space $\mathcal{Y}$ of the type
$\left\|S_{t}\right\|\leq Me^{-\gamma t}\qquad\text{ for }\gamma>0,M>1,t>0$ (24)
(where $\left\|\cdot\right\|$ denotes the operator norm) is equivalent to the property that for every $\lambda>-\gamma$, $\lambda$ is in the resolvent set $\rho(L_{\varepsilon})$ of its infinitesimal generator $(L_{\varepsilon},\mathcal{D}(L_{\varepsilon}))$, and
$\left\|R(\lambda,L_{\varepsilon})^{n}\right\|\leq\frac{M}{(\lambda+\gamma)^{n}}\qquad\text{for every }n\in\mathbb{N},$ (25)
where $R$ denotes the resolvent operator. From Theorem 1.1—in particular from Rmk. 1.2—the condition (24) is satisfied in the Banach space $\mathcal{Y}:=\{f\in L^{1}_{k}:\int_{\mathbb{R}^{d}}f=0\}$. Thus, choosing $n=1$ and $\lambda=0>-\gamma$ in (25), it follows that $L_{\varepsilon}$ has an inverse in $\mathcal{Y}$ with $\left\|L_{\varepsilon}^{-1}\right\|_{\mathcal{Y}}\leq M/\gamma$. This implies that,
$\begin{split}\left\|F_{\varepsilon}-G\right\|_{L^{1}_{k}}&=\left\|F_{\varepsilon}-G\right\|_{\mathcal{Y}}=\left\|L_{\varepsilon}^{-1}L_{\varepsilon}(F_{\varepsilon}-G)\right\|_{\mathcal{Y}}\leq\left\|L_{\varepsilon}^{-1}\right\|_{\mathcal{Y}}\left\|L_{\varepsilon}(F_{\varepsilon}-G)\right\|_{\mathcal{Y}}\\ &\leq\frac{M}{\gamma}\left\|L_{\varepsilon}(F_{\varepsilon}-G)\right\|_{L^{1}_{k}}=\frac{M}{\gamma}\left\|(L_{\varepsilon}-L_{0})G\right\|_{L^{1}_{k}},\end{split}$
since $L_{\varepsilon}F_{\varepsilon}=L_{0}G=0$. Given the regularity and decay properties of the Gaussian, by consistency 4.1 we obtain
$\left\|(L_{\varepsilon}-L_{0})G\right\|_{L^{1}_{k}}\leq C\varepsilon^{2},$
concluding the proof. ∎
As a consequence of the previous proposition and, again, Theorem 1.1 we can estimate the speed of convergence at large times.
Proposition 4.8 (Convergence for large times).
Let $J$ satisfy the hypotheses of Theorem 1.1 $(i)$, Assumption (4) and, additionally $J\in L^{1}_{k+d+4}$. Let $u$ be the solution of (NLFP) and $v$ be the solution of (FP) with the same initial data $u_{0}\in L^{1}_{k}$. Then, there exist $T^{*}>0$ and $C>0$ (depending on $d$, $k$, $p$, and $u_{0}$) such that
$\left\|u(t,\cdot)-v(t,\cdot)\right\|_{L^{1}_{k}}\leq C\varepsilon^{2}\qquad\text{ for all $t>T^{*}$.}$
Proof.
From Theorem 1.1
$\left\|u(t,\cdot)-F_{\varepsilon}\right\|_{L^{1}_{k}}\leq C_{1}e^{-\lambda_{1}t}\left\|u_{0}-F_{\varepsilon}\right\|_{L^{1}_{k}}.$
Moreover, we have a similar result for the local case— see e.g. Gualdani et al. (2018).
$\left\|v(t,\cdot)-G\right\|_{L^{1}_{k}}\leq C_{2}e^{-\lambda_{2}t}\left\|u_{0}-G\right\|_{L^{1}_{k}}$
where $v$ is the solution of the local Fokker Planck equation with the same initial data. This means
$\left\|u(t,\cdot)-v(t,\cdot)\right\|_{L^{1}_{k}}\leq\left\|F_{\varepsilon}-G\right\|_{L^{1}_{k}}+Ce^{-\lambda t}\Big(\left\|u_{0}-F_{\varepsilon}\right\|_{L_{1}^{k}}-\left\|u_{0}-G\right\|_{L_{1}^{k}}\Big).$
From Theorem 4.7,
$\left\|F_{\varepsilon}-G\right\|_{L^{1}_{k}}\leq\tilde{C}\varepsilon^{2}.$
Let us define $T^{*}$ such that
$e^{-\lambda T^{*}}\Big(\left\|u_{0}-F_{\varepsilon}\right\|_{L^{1}_{k}}-\left\|u_{0}-G\right\|_{L^{1}_{k}}\Big)\leq\varepsilon^{2}.$
Then for all $t>T^{*}$
$\left\|u(t,\cdot)-v(t,\cdot)\right\|_{L^{1}_{k}}\leq C\varepsilon^{2}.\qed$
We finally combine the previous results to give the proof of Theorem 1.3:
Proof of Theorem 1.3.
The proof comes directly from the previous estimates in this section. Indeed, take $T^{*}$ as in Proposition 4.8. If $t\in[0,T^{*}]$, then apply Proposition 4.6. If $t>T^{*}$, then Proposition 4.8 gives the result. ∎
5 Analysis of the equilibrium: regularity, moments, cumulants and tails
Thanks to Harris’s Theorem we have proved the existence and uniqueness of an exponentially stable equilibrium for the equation (NLFP) at each $F_{\varepsilon}$. We can infer some information about its regularity and shape, especially on the cumulants, moments and tails.
5.1 Regularity
Thanks to the explicit form (11) of the equilibrium in the frequency domain, one can study the regularity by standard results on the Fourier transform.
Recall that the Fourier transform of the equilibrium is
$\hat{F}_{\varepsilon}(\xi)=\exp\Big(\frac{1}{\varepsilon^{2}}\int_{0}^{\infty}\zeta_{\varepsilon}(e^{-s}\xi)\,\mathrm{d}s\Big)$
with $\zeta_{\varepsilon}(\xi)=\hat{J}_{\varepsilon}(\xi)-1$.
Theorem 5.1.
The equilibrium $F_{\varepsilon}$ of equation (NLFP) has the following regularity properties. For all $\varepsilon\in(0,1]$
  1. (i)
    $F_{\varepsilon}\in\mathcal{C}^{k-d}$ for all $k<\frac{1}{\varepsilon^{2}}$ such that $k\geq d$.
  2. (ii)
    $F_{\varepsilon}\notin\mathcal{C}^{k}$ for all $k>\frac{1}{\varepsilon^{2}}$.
Proof.
Let us assume $\zeta_{\varepsilon}$ radially symmetric. If not, define $\underline{\zeta}_{\varepsilon}$ and $\overline{\zeta}_{\varepsilon}$ by
$\underline{\zeta}_{\varepsilon}(z)=\min_{|y|=|z|}\zeta_{\varepsilon}(y)\qquad\overline{\zeta}_{\varepsilon}(z)=\max_{|y|=|z|}\zeta_{\varepsilon}(y).$
These functions are then radially symmetric and satisfy
$\underline{\zeta}_{\varepsilon}(\xi)\leq\zeta_{\varepsilon}(\xi)\leq\overline{\zeta}_{\varepsilon}(\xi),$
$\underline{\zeta}_{\varepsilon}(\xi)=-|\xi|^{2}+o(|\xi|^{3})\quad\text{and}\quad\overline{\zeta}_{\varepsilon}(\xi)=-|\xi|^{2}+o(|\xi|^{3})\qquad\text{as $\xi\to 0$.}$
After a change of variable $y=e^{-s}|\xi|$, we obtain
$\int_{0}^{\infty}\zeta_{\varepsilon}(e^{-s}|\xi|)\,\mathrm{d}s=\int_{0}^{|\xi|}\frac{\zeta_{\varepsilon}(y)}{y}\,\mathrm{d}y$
Since $\zeta_{\varepsilon}(y)=-\varepsilon^{2}y^{2}+o(y^{2})$, the singularity at zero is canceled and, moreover, close to $\xi=0$ we have
$\hat{F}_{\varepsilon}(\xi)\approx e^{-\frac{|\xi|^{2}}{2}}.$
To study the regularity, we are interested in the decay properties of the Fourier transform
  1. (i)
    To prove the first claim, it suffices to show that
    $|\hat{F}_{\varepsilon}(\xi)|\leq\frac{c}{1+|\xi|^{k+\eta}}\qquad\text{ for }k<\frac{1}{\varepsilon^{2}}$ (26)
    for some (small) positive $\eta$. Indeed, by the Riemann-Lebesgue lemma, $\hat{J}(y)\to 0$ as $|y|\to\infty$. Let $y_{\delta}>0$ be such that
    $|\hat{J}(y)|\leq\delta:=\frac{\varepsilon^{2}}{2}\Big(\frac{1}{\varepsilon^{2}}-k\Big)\qquad\text{ for all }|y|>y_{\delta}$
    Without loss of generality, we can assume $y_{\delta}<|\xi|$, otherwise there is nothing to prove. Then
    $\begin{split}\int_{0}^{|\xi|}\frac{\zeta_{\varepsilon}(y)}{y}\,\mathrm{d}y=\int_{0}^{y_{\delta}}\frac{\zeta_{\varepsilon}(y)}{y}\,\mathrm{d}y+\int_{y_{\delta}}^{|\xi|}\frac{\zeta_{\varepsilon}(y)}{y}\,\mathrm{d}y\leq C_{1}-(1-\delta)(\log|\xi|-\log|y_{\delta}|)\end{split}$
    which implies that |^F_ε(ξ)|≤Cexp(-1-δε2log|ξ|)=C|ξ|1-δε2C|ξ|k+(1ε2-k)/2. For $|\xi|$ large, the previous implies condition (26) with $\eta<(\frac{1}{\varepsilon^{2}}-k)/2$ and by standard argument on Fourier transform $F_{\varepsilon}\in\mathcal{C}^{k-d}$ for all $k<\frac{1}{\varepsilon^{2}}$.
  2. (ii)
    To prove the second claim, we will show that
    $|\xi|^{k}|\hat{F}_{\varepsilon}(\xi)|\notin L^{\infty}\qquad\text{ for }k>\frac{1}{\varepsilon^{2}}$
    proceeding in a similar way as before. Notice that since $\xi\mapsto\hat{J}(\xi)$ is continuous and decays to zero as $|\xi|\to+\infty$, there exists $y_{\delta}$ such that we can bound $\hat{J}$ from below by ^J(x)≥- δ:=ε22(1ε2-k)   for all y≥y_δ. As a consequence we see that
    $\begin{split}|\xi|^{k}\hat{F}_{\varepsilon}(\xi)\geq|\xi|^{k}C\exp\Big(-\frac{1+\delta}{\varepsilon^{2}}\log|x|\Big)=|\xi|^{k-\frac{1+\delta}{\varepsilon^{2}}}\notin L^{\infty}\end{split}$
    and thus by standard properties of the Fourier transform, $F_{\varepsilon}\notin\mathcal{C}^{k}$.
We just proved that the regularity keeps improving when $\varepsilon\to 0$ although the result is not optimal. A precise characterization remains, to our knowledge, an open question. Moreover, we suspect that under quite general conditions, the only point at which $F_{\varepsilon}$ is not $\mathcal{C}^{\infty}$ is the origin. We are able to show this in dimension $1$, and for radially symmetric $J$:
Theorem 5.2.
Let $\Omega\subset\subset\mathbb{R}^{d}$ such that $0\notin\Omega$. Suppose (at least) one of the two following condition is verified
  1. (a)
    $d=1$
  2. (b)
    $J$ is radially symmetric.
then $F_{\varepsilon}\in\mathcal{C}^{\infty}(\mathbb{R}^{d}\setminus\{0\}).$
Proof.
We proceed via a bootstrap argument: the equilibrium satisfies the stationary equation
$\operatorname{div}(xF_{\varepsilon})=-\frac{1}{\varepsilon^{2}}(J_{\varepsilon}*F_{\varepsilon}-F_{\varepsilon})$ (27)
Since $F_{\varepsilon}\in L^{1}(\Omega)$, the RHS is also in $L^{1}(\Omega)$, and thus $\operatorname{div}(xF_{\varepsilon})$ must be in $L^{1}(\Omega)$ as well.
  1. (a)
    If $d=1$ then $\operatorname{div}(xF_{\varepsilon})\in L^{1}(\Omega)$ and $F_{\varepsilon}\in L^{1}(\Omega)$ implies that $xF^{\prime}\in L^{1}(\Omega)$ and, being $\operatorname{dist}(0,\Omega)\geq\delta$, also $F^{\prime}$ is integrable. This, by the fundamental theorem of calculus, leads to continuity for $F$ and consequently for the RHS of (27). Iterating the process one gets the result.
  2. (b)
    If $d\neq 1$, the fact that $\operatorname{div}(xF_{\varepsilon})$ is integrable in $\Omega$ does not assure that $\nabla F\in L^{1}(\Omega)$, even when we are far from 0. Therefore, we need an additional hypothesis. If $J$ is radially symmetric, $F_{\varepsilon}$ inherits this property (see Remark 2.8), and for such functions the following holds: F_ε∈W^j,1(Ω), div(xF_ε)∈W^j,1(Ω) ⟹F_ε∈W^j+1,1(Ω)   for all j≥0 Indeed, since $\operatorname{div}(xF_{\varepsilon})=dF_{\varepsilon}+x\cdot\nabla F_{\varepsilon}$ and, by hypotheses $x\cdot F_{\varepsilon}\in W^{1,1}(\Omega)$, due to radial symmetry we can write x⋅∇F_ε= x⋅F_ε’(|x|)x|x|=F_ε’(|x|)|x|.
[…]
    We then iterate the process as in step $(a)$ using (27), improving the regularity up to $F_{\varepsilon}\in W^{d,1}(\Omega)$, where $d$ is the dimension. By the Sobolev embedding theorem, this implies that $F_{\varepsilon}\in\mathcal{C}(\Omega)$.
    If we keep repeating the argument $(k+1)d$ times, we have that $F_{\varepsilon}\in\mathcal{C}^{k}(\Omega)$ for every $k$ and thus the $\mathcal{C}^{\infty}$ regularity follows.
5.2 Moments and cumulants
Let $\alpha=(\alpha_{1},\dots,\alpha_{d})$ be a multi-index with order $|\alpha|=n$. For a nonnegative Borel measure $f$ on $\mathbb{R}^{d}$ We define the absolute moment of order $n$ by
$M_{n}(f):=\int_{\mathbb{R}^{d}}|x|^{n}f(x)\,\mathrm{d}x,$
(always well-defined but possibly $\infty$) and the $\alpha$-moment by
$m_{\alpha}(f):=\int_{\mathbb{R}^{d}}x^{\alpha}f(x)\,\mathrm{d}x=\int_{\mathbb{R}^{d}}x_{1}^{\alpha_{1}}\cdots x_{d}^{\alpha_{d}}f(x)\,\mathrm{d}x,$
which is well-defined whenever $M_{n}[f]<+\infty$. The moment generating function $\mathfrak{M}(\xi)$ is given by
$\mathfrak{M}[f](\xi):=\int_{\mathbb{R}^{d}}e^{x\xi}f(t,x)\,\mathrm{d}x$
whenever $x\mapsto e^{x\xi}u(t,x)$ is integrable. Moments can then also be defined via the moment generating function by
$m_{\alpha}(f)=D^{\alpha}\mathfrak{M}[f](\xi)\Big|_{\xi=0}.$
Cumulants are an alternative way to describe a distribution, similar to the moments. They can be defined through the cumulative generating function (c.g.f.) $C[f]$, which is the logarithm of the moment generating function $\mathfrak{M}(\xi)$:
$C[f](\xi):=\log\int_{\mathbb{R}^{d}}e^{x\xi}f(x)\,\mathrm{d}x.$
The $\alpha$-cross cumulant is then defined by
$\kappa_{\alpha}(f)=D^{\alpha}C[f](\xi)\Big|_{\xi=0}.$
If two distributions have the same moments then they have the same cumulants and vice versa. However, moments always depend on the moments of lower order while the cumulants don’t.
Using Fáa di Bruno’s formula, a generalization of the chain rule for repetitive derivatives, one can prove that cumulants and moments are linked via Bell’s polynomials, defined in Comtet (1974). These are well known combinatorial objects in $\mathbb{R}$, defined through their exponential generating function, that is
$\exp\left(\sum_{i=1}^{\infty}\frac{x_{i}}{i!}y^{i}\right)=\sum_{n=0}^{\infty}\frac{B_{n}(x_{1},\dots,x_{n})}{n!}y^{n}.$ (28)
In the rest of this section, $u$ denotes a solution to (NLFP), and $C=C(t,\xi)$ is its cumulative generating function:
$C(t,\xi)=\log\int_{\mathbb{R}^{d}}e^{x\xi}u(t,x)\,\mathrm{d}x.$
This can be explicitly computed in a simple way:
Theorem 5.3.
Let $J$ satisfy (1) and let $u$ be the solution of (NLFP) with initial data $u_{0}$. Assume that $\int_{\mathbb{R}^{d}}e^{|x|}u_{0}\,\mathrm{d}x<\infty$ and $\int_{\mathbb{R}^{d}}e^{|x|}J\,\mathrm{d}x<\infty$. Then
$C(t,\xi)=C_{0}(e^{-t}\xi)+\frac{1}{\varepsilon^{2}}\int_{0}^{t}\zeta(e^{-s}\xi)\,\mathrm{d}s$
with $C_{0}(\xi)=\log\int_{\mathbb{R}^{d}}e^{x\xi}u_{0}(x)\,\mathrm{d}x$ and $\zeta(\xi)=\mathfrak{M}(J)-1=\int_{\mathbb{R}^{d}}e^{x\xi}J(x)\,\mathrm{d}x-1$.
Proof.
We proceed similarly to the work of Alfaro and Veruete (2020) and Gil et al. (2017). The partial derivatives of the c.g.f. are
$\partial_{\xi_{i}}C(t,\xi)=\frac{\int_{\mathbb{R}^{d}}x_{i}e^{x\cdot\xi}\,\mathrm{d}x}{\int_{\mathbb{R}^{d}}e^{x\cdot\xi}\,\mathrm{d}x}$
and
$\partial_{t}C(t,\xi)=\frac{\int_{\mathbb{R}^{d}}e^{x\xi}\partial_{t}u\,\mathrm{d}x}{\int_{\mathbb{R}^{d}}e^{x\xi}u\,\mathrm{d}x}=\frac{1}{\varepsilon^{2}}\frac{\int_{\mathbb{R}^{d}}e^{x\xi}(J_{\varepsilon}*u-u)\,\mathrm{d}x}{\int_{\mathbb{R}^{d}}e^{x\xi}u\,\mathrm{d}x}+\frac{\int_{\mathbb{R}^{d}}e^{x\xi}\operatorname{div}(xu)\,\mathrm{d}x}{\int_{\mathbb{R}^{d}}e^{x\xi}u\,\mathrm{d}x}$
Analyzing the first term we get
$\begin{split}&\int_{\mathbb{R}^{d}}e^{x\xi}(J_{\varepsilon}*u)\,\mathrm{d}x=\int_{\mathbb{R}^{d}}e^{x\xi}\Big(\int_{\mathbb{R}^{d}}J_{\varepsilon}(x-y)u(t,y)\,\mathrm{d}y\Big)\,\mathrm{d}x\\ &=\int_{\mathbb{R}^{d}}e^{y\xi}u(t,y)\,\mathrm{d}y\int_{\mathbb{R}^{d}}J_{\varepsilon}(x-y)e^{(x-y)\xi}\,\mathrm{d}x=\Big(\int_{\mathbb{R}^{d}}e^{y\xi}u(t,y)\,\mathrm{d}y\Big)\mathfrak{M}(J_{\varepsilon}).\end{split}$
For the second term, integrating by parts,
$\int_{\mathbb{R}^{d}}e^{x\cdot\xi}\operatorname{div}_{x}(xu)\,\mathrm{d}x=-\int_{\mathbb{R}^{d}}\nabla_{x}(e^{x\cdot\xi})\cdot(xu)=-\xi\cdot\int_{\mathbb{R}^{d}}e^{x\cdot\xi}xu=-\xi\cdot\nabla_{\xi}C(t,\xi)\int_{\mathbb{R}^{d}}e^{x\xi}u\,\mathrm{d}x$
Therefore
$\partial_{t}C(t,\xi)=\frac{1}{\varepsilon^{2}}\zeta_{\varepsilon}(\xi)-\xi\cdot\nabla_{\xi}C(t,\xi).$
One can check that the only solution to this above first order PDE is the function $C$ in the statement (of course, the PDE can be solved, for example by the method of characteristics, to obtain the given formula). Hence the proof is complete. ∎
Corollary 5.4.
In the hypotheses of the previous theorem, let $\alpha$ be a multi-index with $|\alpha|=n$. Then,
  1. (i)
    The $\alpha$-cross cumulant is
    $\kappa_{\alpha}(u(t))=e^{-|\alpha|t}\kappa_{\alpha}(u_{0})+(1-e^{-|\alpha|t})\frac{1}{|\alpha|\varepsilon^{2}}m_{\alpha}(J_{\varepsilon})$
    As a consequence
    $\kappa_{\alpha}(F_{\varepsilon})=\frac{1}{|\alpha|\varepsilon^{2}}m_{\alpha}(J_{\varepsilon})$
  2. (ii)
    If $d=1$, then, the nthn-th moment is
    $m_{n}(u)=B_{n}(\kappa_{1},\dots,\kappa_{n})=\sum_{k=1}^{n}B_{n,k}(\kappa_{1},\dots,\kappa_{n-k+1})$ (29)
    where $B_{n}$ are the complete (exponential) Bell’s Polynomials defined in (28) and $B_{n,k}$ are called partial Bell’s Polynomials [^1] From a combinatorial point of view, $B_{n,k}(x_{1},\dots,x_{n-k+1}$) enumerates the possible ways to partition a set of $n$ elements into $k$ non-empty subsets. Each term of the polynomials represents specific way of distributing the $n$ elements in the $k$ subsets. For example, $B_{5,2}=10x_{2}x_{3}+5x_{1}x_{4}$ means that there are $10$ ways to partition a set of $5$ elements into two subsets of cardinality respectively $2$ and $3$ and $5$ ways to do so in two subsets of cardinality $1$ and $4$. .
Proof.
From Theorem 5.3 the $\alpha$-th derivative of the c.g.f. is
$D^{\alpha}C(t,\xi)=e^{-nt}(D^{\alpha}C_{0})(e^{-t}\xi)+\frac{1-e^{-nt}}{n\varepsilon^{2}}(D^{\alpha}\beta)(e^{-t}\xi)$
and thus, evaluating at $\xi=0$ we get the $\alpha$-mixed cumulant
$\kappa_{\alpha}(u(t))=e^{-nt}\kappa_{\alpha}(u_{0})+(1-e^{-nt})\frac{1}{n\varepsilon^{2}}m_{\alpha}(J_{\varepsilon})$
Expression (29) comes directly from the definition of Bell’s polynomials (28). ∎
Remark 5.5.
One can generalize the last result to higher dimension but it appears that the multidimensional version of Bell’s polynomials is not very well known. We refer to Withers and Nadarajah (2010), Schumann (2019) and references therein for further information.
Remark 5.6.
By computing explicitly the time derivative of the $\alpha$ moment, i.e.
$\begin{split}\frac{\mathrm{d}}{\mathrm{d}t}m_{\alpha}(u(t))=\frac{1}{\varepsilon^{2}}(m_{\alpha}(J_{\varepsilon}*u)-m_{\alpha}(u))-m_{\alpha}(\operatorname{div}(xu)),\end{split}$
one can obtain the following recursive formula for the moments of the solution (and consequently for the equilibrium):
$m_{\alpha}(u(t))=\gamma_{\alpha}(1-e^{-nt})+e^{-nt}m_{\alpha}(u_{0})$
where
$\gamma_{\alpha}=\frac{1}{n\varepsilon^{2}}\sum_{\begin{subarray}{c}\beta\leq\alpha\\ \beta\neq 0\end{subarray}}^{n}\binom{\alpha}{\beta}m_{\beta}(J)m_{\alpha-\beta}(u).$
5.3 Tails
For the usual (local) equilibrium for the Fokker-Planck equation (FP), is a standard Gaussian. In the nonlocal approximation in this paper, the tails of the equilibrium seem to decay slower, no matter how fast the kernel $J$ is decaying. As a consequence of Harris’s theorem and our Lyapunov condition estimates (Theorem 1.3 (ii)), we know for an exponentially decaying kernel $J\in L_{\exp,a^{\prime}}$, the unique probability equilibrium has exponentially decaying tails; that is, for every $a<a^{\prime}\varepsilon^{-1}$
$\int_{\mathbb{R}^{d}}F_{\varepsilon}(x)e^{a\left\langle x\right\rangle}\,\mathrm{d}x<\infty.$
We conjecture that for a kernel $J$ decaying fast enough, the tails of the equilibrium have a “Poisson-like” behaviour: they decay roughly like $e^{-|x|\log(|x|+1)}$. This implies that
$\int_{\mathbb{R}^{d}}F_{\varepsilon}(x)e^{a\left\langle x\right\rangle^{\gamma}}\,\mathrm{d}x=\infty\quad\text{ for every }\gamma>1.$
We have not been able to prove this “Poisson-like” behavior in general, but we can give a good upper bound when $J$ has compact support:
Proposition 5.7.
Let $J:\mathbb{R}^{d}\to\mathbb{R}$ be compactly supported and its support is in the ball $B_{R}$ for some $R>0.$ Then
$\int F_{\varepsilon}(x)e^{a\left\langle x\right\rangle\log\left\langle x\right\rangle}\,\mathrm{d}x<\infty$
for any $a\leq\frac{1}{\varepsilon R}$.
Proof.
Let $\varphi(x)=\exp(a\left\langle x\right\rangle\log\left\langle x\right\rangle)$ with $a$ to be determined later. In the same spirit of the proofs of theorems 3.2 and 3.4, one can compute
$\begin{split}&\frac{1}{\varepsilon^{2}}\int_{\mathbb{R}^{d}}J_{\varepsilon}(y)(\varphi(x+y)-\varphi(x))\,\mathrm{d}y=\frac{1}{\varepsilon^{2}}\int_{\mathbb{R}^{d}}J(z)(\varphi(x+\varepsilon z)-\varphi(x))\,\mathrm{d}z\\ &=\frac{1}{\varepsilon^{2}}\int_{\mathbb{R}^{d}}J(z)\Big(\varphi(x)+\frac{1}{2}\sum_{|\alpha|=2}R_{\alpha}(\xi)\varepsilon^{2}z^{\alpha}-\varphi(x)\Big)\,\mathrm{d}z=\frac{1}{2}\int_{\mathbb{R}^{d}}J(z)\sum_{i=1}^{d}\partial_{i}^{2}\varphi(\xi)z_{i}^{2}\,\mathrm{d}z.\end{split}$
With a simple calculation
$\begin{split}\sum_{i=1}^{d}\partial_{i}^{2}\varphi(x)&=\sum_{i=1}^{d}a\frac{\varphi}{\left\langle x\right\rangle^{3}}\Big(a^{2}x_{i}^{2}\left\langle x\right\rangle(\log^{2}\left\langle x\right\rangle+2\log\left\langle x\right\rangle+1)+(1+\sum_{j\neq 1}x_{j}^{2})(\log\left\langle x\right\rangle+1)\Big)\\ &\leq C_{1}a^{2}\log^{2}\left\langle x\right\rangle\varphi(x),\end{split}$
and since $\varphi$ is radially increasing $\left\langle|x|+|z|\right\rangle\leq\left\langle x\right\rangle+\left\langle z\right\rangle$, and $\log^{2}\left\langle x+z\right\rangle\lesssim\log^{2}{\left\langle x\right\rangle}+\log^{2}\left\langle z\right\rangle$ we have
$\begin{split}\frac{1}{\varepsilon^{2}}&\int_{\mathbb{R}^{d}}J_{\varepsilon}(y)(\varphi(x+y)-\varphi(x))\,\mathrm{d}y\leq\frac{C_{1}}{2}a^{2}\int_{B_{R}}J(z)\log^{2}\left\langle\xi\right\rangle\varphi(\xi)\,\mathrm{d}z\\ &\leq C_{2}a^{2}\log^{2}\left\langle x\right\rangle\varphi(|x|+\varepsilon R)\int_{B_{R}}J(z)\,\mathrm{d}z=C_{2}a^{2}\log^{2}\left\langle x\right\rangle\left\langle|x|+\varepsilon R\right\rangle^{a\left\langle|x|+\varepsilon R\right\rangle}\\ &\leq C_{2}a^{2}\log^{2}\left\langle x\right\rangle\left\langle|x|+\varepsilon R\right\rangle^{a\varepsilon R}e^{a\left\langle x\right\rangle\log\left\langle x\right\rangle}=\Big[C_{2}a^{2}\log^{2}\left\langle x\right\rangle\left\langle|x|+\varepsilon R\right\rangle^{a\varepsilon R}\Big]\varphi(x)\end{split}$
Moreover,
$\begin{split}x\cdot\nabla\varphi&=x\cdot a\frac{x}{\left\langle x\right\rangle}(\log\left\langle x\right\rangle+1)\varphi(x)\leq\Big[C_{3}a\left\langle x\right\rangle\log\left\langle x\right\rangle\Big]\varphi(x)\end{split}$
Consequently,
$\begin{split}L^{*}_{\varepsilon}&\varphi(x)=\frac{1}{\varepsilon^{2}}\int_{\mathbb{R}^{d}}J_{\varepsilon}(y)(\varphi(x+y)-\varphi(x))\,\mathrm{d}y-x\cdot\nabla\varphi(x)\\ &\leq\Big[C_{2}a^{2}\log^{2}\left\langle x\right\rangle\left\langle|x|+\varepsilon R\right\rangle^{a\varepsilon R}-C_{3}a\left\langle x\right\rangle\log\left\langle x\right\rangle\Big]\varphi(x)\leq C-\lambda\varphi\end{split}$ (30)
for some $C$ and $\lambda>0$ and $a<\frac{1}{\varepsilon R}$. ∎
Acknowledgments
We would like to thank Josephine Evans for discussions which gave rise to the early ideas of this paper. We would also like to thank Alejandro Gárriz for several discussions and suggestions.
The project that gave rise to these results received the support of a fellowship from the “la Caixa” Foundation (ID 100010434). The fellowship code is LCF/BQ/DI22/11940032. Both authors acknowledge support from the “Maria de Maeztu” Excellence Unit IMAG, reference CEX2020-001105-M, funded by MCIN/AEI/10.13039/501100011033/. They were also supported by grant PID2023-151625NB-I00 and the research network RED2022-134784-T from MCIN/AEI/10.13039/501100011033/.